### v1

# Reactive capacitance of flat patches of arbitrary shape

Abstract
We investigate the capacity of a flat partially reactive patch of arbitrary shape to trap independent particles that undergo steady-state diffusion in the three-dimensional space. We focus on the total flux of particles onto the patch that determines its reactive capacitance. To disentangle the respective roles of the reactivity and the shape of the patch, we employ a spectral expansion of the reactive capacitance over a suitable Steklov eigenvalue problem. We derive several bounds on the reactive capacitance to reveal its monotonicity with respect to the reactivity and the shape. Two probabilistic interpretations are presented as well. An efficient numerical tool is developed for solving the associated Steklov spectral problem for patches of arbitrary shape. We propose and validate, both theoretically and numerically, a simple, fully explicit approximation for the reactive capacitance that depends only on the surface area and the electrostatic capacitance of the patch. This approximation opens promising ways to access various characteristics of diffusion-controlled reactions in general domains with multiple small well-separated patches. Direct applications of these results in statistical physics and physical chemistry are discussed.
diffusion, capacitance, Steklov problem, partial reactivity, Robin boundary condition, first-reaction time
pacs:
02.50.-r, 05.40.-a, 02.70.Rr, 05.10.Gg
I Introduction
In many biochemical scenarios, a particle diffuses in a complex medium and searches for small reactive patches that are dispersed on reflecting walls [1, 2, 3, 4, 5]. Depending on the considered application, the reactive patch can represent an ion channel, an active site or a catalytic germ, a hole, a magnetic impurity, etc. The mean first-passage time (MFPT) to the patch is often used to characterize the efficiency of the diffusive search [6, 7, 8, 9, 10]. Starting from the Lord Rayleigh’s seminal result for the MFPT to a small circular patch on the sphere [11], considerable progress has been achieved in solving this so-called narrow escape problem [12, 13, 14, 15, 16, 17, 18, 19, 20, 21] (see also overviews in [22, 23]). Among various analytical tools, one can mention the method of matched asymptotic expansions [24, 25, 26, 27, 28, 29, 30, 31], homogenization techniques [32, 33, 34, 35, 36, 37], constant-flux approximation [38, 39, 40, 41, 42, 43], Wiener-Hopf integral equation [44], as well as conformal mapping in two dimensions [45, 46]. Other characteristics of diffusion-controlled reactions [47, 48, 49, 50, 51, 52] such as the overall reaction rate (or the total flux) and splitting probabilities were also analyzed.
Former works were mainly focused on the idealized situation of perfectly reactive patches when the reaction event occurs immediately upon the first contact with the patch. In other words, the kinetic step of the reaction event was ignored, by assuming that it occurs much faster than the diffusion step. However, in most practical settings, once the particle has arrived onto the patch, it has to overpass an energy activation barrier to react, or an entropic barrier to squeeze through a hole. In biochemistry, a macromolecule has to be in the proper conformational state to bind its partner, an ion channel has to be open to pass the arrived ion, and so on. As discussed in [39], the ignorance of this kinetic step may cause misleading conclusions on the asymptotic behavior of the mean first-reaction time (MFRT) and other characteristics. The important role of partial reactivity in diffusion-controlled reactions, first recognized in the seminal work by Collins and Kimball [53], has been revealed in different contexts [54, 55, 56, 57, 58, 59, 60, 61, 64, 62, 63]. In particular, the asymptotic behavior of the MFRT on partially reactive targets was shown to be considerably different from the idealized case of perfect reactions [64, 65, 66]. Recently, the problem of diffusion-controlled reactions on multiple small well-separated partially reactive patches in both two and three dimensions was revisited [68, 67, 69]. In particular, the MFRT in three dimensions was shown to be determined, to the leading order, by the reactive capacitance, which is proportional to the flux of particles reacted on the patch and thus naturally generalizes the electrostatic capacitance to the case of partial reactivity (see Sec. II for details). For a circular patch, the reactive capacitance was thoroughly investigated in [68].
In this paper, we push this analysis to another direction and investigate the reactive capacitance for patches of various shapes. First, we deepen the theoretical description of the reactive capacitance in Sec. II by uncovering its monotonicity properties and probabilistic interpretations. We also get useful bounds on the reactive capacitance and the principal Steklov eigenvalue. For this purpose, we reformulate the exterior Steklov problem for flat patches as an eigenvalue problem for an integral operator with an explicit kernel. Second, we employ this reformulation to conceive an efficient method for solving the exterior Steklov problem numerically. Various numerical results are present in Sec. III. In particular, we inspect how the Steklov eigenvalues and eigenfunctions depend on the shape of the patch by considering several families of shapes such as ellipses and rectangles. We show that the principal eigenfunction provides the dominant contribution to the reactive capacitance and validate a simple, fully explicit approximation for this quantity. Section IV discusses physical implications of this analysis and concludes the paper by summarizing the main results and highlighting open questions.
Figure 1: (a) A domain $\Omega$ with reflecting boundary $\partial\Omega_{0}$ (in gray) A flat patch $\Gamma$ on the horizontal plane of the upper half-space.
II Reactive capacitance
Throughout this paper, we consider a flat partially reactive bounded patch $\Gamma$ of arbitrary shape on the horizontal reflecting plane in the upper half-space ${\mathbb{R}}^{3}_{+}$ (Fig. 1). To avoid eventual mathematical issues on the well-posedness of the considered problems, we assume that the boundary of the patch is piecewise smooth. We consider steady-state diffusion of particles with constant diffusion coefficient $D>0$, concentration $A_{0}>0$ imposed at infinity, and reactivity $\kappa>0$ on the patch. The concentration profile in the upper half-space can be expressed as $A(\bm{y})=A_{0}(1-w(\bm{y};\mu))$, where $\mu=\kappa/D$ and the function $w(\bm{y};\mu)$ solves the boundary value problem
$\Delta w$ $=0\quad(\bm{y}\in{\mathbb{R}}^{3}_{+}),$ (1a)
$\partial_{n}w+\mu w$ $=\mu\quad(\bm{y}\in\Gamma),$ (1b)
$\partial_{n}w$ $=0\quad(\bm{y}\in\partial{\mathbb{R}}^{3}_{+}\backslash\overline{\Gamma}),$ (1c)
$w$ $\sim\frac{C(\mu)}{|\bm{y}|}+\mathcal{O}(|\bm{y}|^{-2})\quad(|\bm{y}|\to\infty),$ (1d)
where $\bm{y}=(y_{1},y_{2},y_{3})$ and $\partial_{n}=-\partial_{y_{3}}$ is the normal derivative. The coefficient $C(\mu)$ in the far-field behavior of the solution is called the reactive capacitance of the patch $\Gamma$. More explicitly, the divergence theorem yields
$C(\mu)=\frac{1}{2\pi}\int\limits_{\Gamma}(\partial_{n}w)d\bm{y},$ (2)
i.e., $C(\mu)$ is proportional to the flux of particles onto a partially reactive patch $\Gamma$. In the limit $\mu\to\infty$, one retrieves the electrostatic capacitance (also known as harmonic or Newtonian capacity) of the patch $\Gamma$ [70], as if it was perfectly reactive (absorbing). Note that the solution $w$ can be extended to the lower half-space by symmetry so that the considered setting is actually equivalent to a “two-sided” patch in the whole space, without the auxiliary Neumann condition (1c). However, as the patch is flat (of zero thickness), it is more convenient to deal with the problem (1) in the upper half-space.
Even for a circular patch, there is no explicit solution of the boundary value problem (1). In order to investigate the dependence of the reactive capacitance on the reactivity parameter $\mu$ and the patch shape, we will employ spectral expansions, variational formulations, probabilistic interpretations and numerical analysis.
II.1 Exterior Steklov problem
In order to obtain a spectral expansion of the reactive capacitance, we consider the exterior Steklov problem in the upper half-space ${\mathbb{R}}^{3}_{+}$, which consists in finding the eigenpairs $\{\mu_{k},\Psi_{k}\}$ that satisfy [68]
$\Delta\Psi_{k}$ $=0\quad(\bm{y}\in{\mathbb{R}}^{3}_{+}),$ (3a)
$\partial_{n}\Psi_{k}$ $=\mu_{k}\Psi_{k}\quad(\bm{y}\in\Gamma),$ (3b)
$\partial_{n}\Psi_{k}$ $=0\quad(\bm{y}\in\partial{\mathbb{R}}^{3}_{+}\backslash\overline{\Gamma}),$ (3c)
$\Psi_{k}$ $\to 0\quad(|\bm{y}|\to\infty).$ (3d)
This is a special case of the Steklov eigenvalue problem for the exterior of a compact domain [71, 72, 73], which is known to have a discrete spectrum, i.e., a countable set of eigenvalues $\mu_{k}$ that can be enumerated by an integer index $k=0,1,\ldots$ in an increasing order:
$0<\mu_{0}\leq\mu_{1}\leq\ldots\nearrow+\infty.$ (4)
The first eigenvalue $\mu_{0}$ is simple and strictly positive, whereas the associated eigenfunction $\Psi_{0}$ does not change sign. The restrictions of the Steklov eigenfunctions onto the patch $\Gamma$, $\Psi_{k}|_{\Gamma}$, form a complete orthonormal basis of the $L^{2}(\Gamma)$ space of square-integrable functions on $\Gamma$:
$\int\limits_{\Gamma}\Psi_{k}(\bm{y})\,\Psi_{k^{\prime}}(\bm{y})d\bm{y}=\delta_{k,k^{\prime}}.$ (5)
(see [74, 75, 76] for a broad overview of spectral properties of the conventional interior Steklov problem).
To gain a more practical insight onto this spectral problem, let us introduce the Green’s function $G(\bm{y},\bm{y}^{\prime})$ in the upper half-space ${\mathbb{R}}^{3}_{+}$ with Neumann boundary condition on the horizontal plane $\partial{\mathbb{R}}^{3}_{+}$:
$-\Delta G(\bm{y},\bm{y}^{\prime})$ $=\delta(\bm{y}-\bm{y}^{\prime})\quad(\bm{y}\in{\mathbb{R}}^{3}_{+}),$ (6a)
$\partial_{n}G(\bm{y},\bm{y}^{\prime})$ $=0\quad(\bm{y}\in\partial{\mathbb{R}}^{3}_{+}),$ (6b)
$G(\bm{y},\bm{y}^{\prime})$ $\to 0\quad(|\bm{y}|\to\infty),$ (6c)
where $\delta(\bm{y}-\bm{y}^{\prime})$ is the Dirac distribution. By the method of images, one gets immediately
$G(\bm{y},\bm{y}^{\prime})=\frac{1}{4\pi}\biggl(\frac{1}{|\bm{y}-\bm{y}^{\prime}|}+\frac{1}{|\bm{y}-\bar{\bm{y}}^{\prime}|}\biggr),$ (7)
where $\bar{\bm{y}}^{\prime}$ is the mirror reflection of $\bm{y}^{\prime}$ with respect to the horizontal plane. Multiplying Eq. (3a) by $G(\bm{y},\bm{y}^{\prime})$, multiplying Eq. (6a) by $\Psi_{k}(\bm{y})$, subtracting these equations, integrating them over $\bm{y}\in{\mathbb{R}}^{3}_{+}$, using the Green’s formula and the boundary conditions, one finds
$\Psi_{k}(\bm{y}^{\prime})=\mu_{k}\int\limits_{\Gamma}G(\bm{y},\bm{y}^{\prime})\Psi_{k}(\bm{y})d\bm{y}\quad(\bm{y}^{\prime}\in\overline{{\mathbb{R}}^{3}_{+}}).$ (8)
Restricting $\bm{y}^{\prime}$ to $\Gamma$, one can introduce an integral operator $\mathcal{G}$, acting on a continuous function $f(\bm{y})$ on $\Gamma$ as
$(\mathcal{G}f)(\bm{y})=\int\limits_{\Gamma}\mathcal{G}(\bm{y},\bm{y}^{\prime})f(\bm{y}^{\prime})d\bm{y}^{\prime},$ (9)
with the kernel
$\mathcal{G}(\bm{y},\bm{y}^{\prime})=\frac{1}{2\pi|\bm{y}-\bm{y}^{\prime}|}\qquad(\bm{y},\bm{y}^{\prime}\in\Gamma,~\bm{y}\neq\bm{y}^{\prime})$ (10)
(even though the operator $\mathcal{G}$ is formally defined on the space of continuous functions, it can be extended to $L^{2}(\Gamma)$ space via embedding theorems, see [77] for details). Since the kernel $\mathcal{G}(\bm{y},\bm{y}^{\prime})$ is symmetric and weakly singular, $\mathcal{G}$ is a self-adjoint compact operator [77]; moreover, if $\Gamma$ was replaced by the whole plane ${\mathbb{R}}^{2}$, $\mathcal{G}$ would be the Riesz potential of order $1$ [78], which is known to be positive definite (this is a direct corollary of the Fourier transform representation of Riesz potentials, see [79]). As a consequence, there is an infinite sequence of positive eigenvalues, enumerated by the index $k=0,1,2,\ldots$ that accumulate to 0, whereas the associated eigenfunctions form an orthonormal basis of $L^{2}(\Gamma)$. Rewriting Eq. (8) as
$\mathcal{G}\Psi_{k}|_{\Gamma}=\frac{1}{\mu_{k}}\Psi_{k}|_{\Gamma}\qquad(k=0,1,2,\ldots),$ (11)
one sees that $1/\mu_{k}$ and $\Psi_{k}|_{\Gamma}$ are the eigenvalues and eigenfunctions of $\mathcal{G}$. This is an alternative way to access the eigenvalues and eigenfunctions of the Steklov problem (3); note that once $\Psi_{k}|_{\Gamma}$ is found, its extension to the upper half-space follows immediately from Eq. (8). We finally stress that the integral operator $\mathcal{G}$ is the inverse of the Dirichlet-to-Neumann operator associated to the exterior Steklov problem (3).
This reformulation reduces the original three-dimensional problem in the unbounded domain (the upper half-space) to an equivalent eigenvalue problem for the integral operator $\mathcal{G}$ with the explicit kernel on the two-dimensional bounded patch $\Gamma$. Note that a modified integral equation with the kernel (10), in which the right-hand side is a constant, was employed by Bernoff and Lindsay to investigate diffusive capture rates for an array of perfectly reactive circular patches [30]. The eigenvalue problem (11) will serve us to develop an efficient numerical tool for computing the Steklov eigenfunctions (see Appendix A). Moreover, it yields a variational (minimax) characterization of the Steklov eigenvalues [77]
$\frac{1}{\mu_{0}}$ $=\sup\limits_{\|u\|=1}\bigl\{(\mathcal{G}u,u)_{L^{2}(\Gamma)}\bigr\},$ (12a)
$\frac{1}{\mu_{k}}$ $=\inf\limits_{u_{1},\ldots,u_{k}}\sup\limits_{u\perp\{u_{1},\ldots,u_{k}\}\atop\|u\|=1}\bigl\{(\mathcal{G}u,u)_{L^{2}(\Gamma)}\bigr\}\quad(k=1,2\ldots),$ (12b)
for test functions $u$ and $u_{k}$ from $L^{2}(\Gamma)$ of the unit $L^{2}(\Gamma)$-norm. This variational form implies the domain monotonicity property:
$\Gamma_{1}\subset\Gamma_{2}\quad\Rightarrow\quad\mu_{k}^{\Gamma_{1}}\geq\mu_{k}^{\Gamma_{2}}\quad(k=0,1,\ldots).$ (13)
This property can be shown by a standard construction, namely, by using a test function $u(\bm{y})$, which is equal to the $k$-th eigenfunction $\Psi_{k}^{\Gamma_{1}}(\bm{y})$ for $\bm{y}\in\Gamma_{1}$ and 0 otherwise:
$\frac{1}{\mu_{k}^{\Gamma_{1}}}=\bigl(\mathcal{G}\Psi_{k}^{\Gamma_{1}}|_{\Gamma_{1}},\Psi_{k}^{\Gamma_{1}}|_{\Gamma_{1}}\bigr)_{L^{2}(\Gamma_{1})}=\bigl(\mathcal{G}u,u\bigr)_{L^{2}(\Gamma_{2})}\leq\frac{1}{\mu_{k}^{\Gamma_{2}}}\,$,
where the last inequality follows from the minimax characterization (12b).
We emphasize that the domain monotonicity property does not hold for the Steklov problem in a more general setting of the exterior of a compact set. For instance, if $\Gamma$ is an oblate spheroidal surface with semi-axes $a,a,b$, the domain monotonicity property (13) fails, as illustrated in Fig. 7 of [80], which shows the non-monotonous dependence of the principal Steklov eigenvalue on the aspect ratio $a/b$. This example highlights that the domain monotonicity property for patches strongly relies on the flatness of these domains.
II.2 Spectral expansion and basic properties
Since the Steklov eigenfunctions $\Psi_{k}|_{\Gamma}$ form the orthonormal basis of $L^{2}(\Gamma)$, they can be used to expand $w(\bm{y};\mu)|_{\Gamma}$ and thus to determine the normal derivative of $w(\bm{y};\mu)$. Substituting this expansion into Eq. (2), one gets the spectral representation of the reactive capacitance (see [68, 81] for details)
$C(\mu)=\frac{\mu|\Gamma|}{2\pi}\sum\limits_{k=0}^{\infty}\frac{\mu_{k}F_{k}}{\mu_{k}+\mu}\,,\quad F_{k}=\frac{1}{|\Gamma|}\biggl(\int\limits_{\Gamma}\Psi_{k}(\bm{y})d\bm{y}\biggr)^{2}.$ (14)
By expanding a constant function on the basis of eigenfunctions $\Psi_{k}|_{\Gamma}$, one can easily check that
$\sum\limits_{k=0}^{\infty}F_{k}=1,$ (15)
so that the coefficient $F_{k}\geq 0$ of the spectral expansion (14) can be interpreted as the relative contribution of the $k$-th Steklov eigenfunction to the reactive capacitance $C(\mu)$. In particular, the $F_{0}$ characterizes the closeness of the principal eigenfunction $\Psi_{0}$ to a constant. A similar spectral expansion was earlier derived for the spectroscopic impedance of an electrochemical cell, though in a slightly different mathematical setting [82].
Evaluating the derivative of $C(\mu)$,
$\frac{dC(\mu)}{d\mu}=\frac{|\Gamma|}{2\pi}\sum\limits_{k=0}^{\infty}\frac{\mu_{k}^{2}F_{k}}{(\mu_{k}+\mu)^{2}}>0,$ (16)
one sees that $C(\mu)$ is a monotonously growing function of $\mu$ for all $\mu\geq 0$. A Taylor expansion of $C(\mu)$ around 0 reads
$C(\mu)=-\sum\limits_{n=1}^{\infty}c_{n}(-\mu)^{n},\quad\textrm{with}~c_{n}=\frac{|\Gamma|}{2\pi}\sum\limits_{k=0}^{\infty}\frac{F_{k}}{\mu_{k}^{n-1}}\,.$ (17)
In [68], alternative representations for the first three coefficients were derived:
$c_{1}=\frac{|\Gamma|}{2\pi},~~c_{2}=\frac{1}{2\pi}\int\limits_{\Gamma}\omega_{\Gamma}(\bm{y})\,d\bm{y},~~c_{3}=\frac{1}{2\pi}\int\limits_{\Gamma}\omega_{\Gamma}^{2}(\bm{y})\,d\bm{y},$ (18)
where
$\omega_{\Gamma}(\bm{y})=\int\limits_{\Gamma}\frac{d\bm{y}^{\prime}}{2\pi|\bm{y}-\bm{y}^{\prime}|}\,.$ (19)
In the following, we will also employ the notation:
$\mathcal{A}_{\Gamma}=\frac{1}{|\Gamma|^{2}}\int\limits_{\Gamma}\omega_{\Gamma}(\bm{y})d\bm{y}=\frac{1}{|\Gamma|^{2}}\int\limits_{\Gamma\times\Gamma}\mathcal{G}(\bm{y},\bm{y}^{\prime})d\bm{y}d\bm{y}^{\prime}.$ (20)
This constant, re-emerging in various contexts [66], will play an important role in the analysis of the reactive capacitance; note that $c_{2}=\mathcal{A}_{\Gamma}|\Gamma|^{2}/(2\pi)$.
In the opposite limit of large $\mu$, Eq. (14) yields the spectral expansion for the electrostatic capacitance:
$C(\infty)=\frac{|\Gamma|}{2\pi}\sum\limits_{k=0}^{\infty}\mu_{k}F_{k}$ (21)
(here we employ the convention that the capacitance of a ball of radius $R$ is equal to $R$).
If the patch is dilated by a factor $\alpha$, i.e., $\Gamma^{\prime}=\alpha\Gamma$, the Steklov eigenvalues are rescaled by $1/\alpha$, $\mu_{k}^{\prime}=\mu_{k}/\alpha$, $F_{k}$ remain unchanged, so that the reactive capacitance of the rescaled patch reads
$C^{\prime}(\mu)=\alpha\,C(\alpha\mu).$ (22)
II.3 Alternative spectral expansion
The reactive capacitance $C(\mu)$ admits an alternative spectral expansion, which is based on the eigenpairs $\{\mu_{k}^{N},\Psi_{k}^{N}(\bm{y})\}$ of the following exterior Steklov problem:
$\Delta\Psi_{k}^{N}$ $=0\quad(\bm{y}\in{\mathbb{R}}^{3}_{+}),$ (23a)
$\partial_{n}\Psi_{k}^{N}$ $=\mu_{k}^{N}\Psi_{k}^{N}\quad(\bm{y}\in\Gamma),$ (23b)
$\partial_{n}\Psi_{k}^{N}$ $=0\quad(\bm{y}\in\partial{\mathbb{R}}^{3}_{+}\backslash\overline{\Gamma}),$ (23c)
$|\bm{y}|^{2}|\nabla\Psi_{k}^{N}|$ $\to 0\quad(|\bm{y}|\to\infty).$ (23d)
While the decay condition (3d) could be interpreted as a Dirichlet condition at infinity, the decay of the gradient of an eigenfunction in Eq. (23d) can be understood as a Neumann condition at infinity (that is reflected by the superscript ’N’). The latter condition was studied in [83, 87, 84, 85, 86] in the context of sloshing or ice-fishing problems in hydrodynamics (see also [88, 89] and references therein); moreover, the eigenpairs $\{\mu_{k}^{N},\Psi_{k}^{N}\}$ naturally appear in the asymptotic analysis of the mixed Steklov-Neumann problem [66, 68]. In Appendix A, we present an efficient numerical tool to solve the spectral problem (23).
The distinction between the spectral problems (3) and (23) becomes clearer from the form of their asymptotic decay at infinity. For instance, in the exterior of a ball, a general isotropic solution of the Laplace equation is a linear combination $c_{N}+c_{D}/|\bm{y}|$ with arbitrary constants $c_{N}$ and $c_{D}$. The Dirichlet condition (3d) imposes $c_{N}=0$, whereas the Neumann condition (23d) imposes $c_{D}=0$. For general domains (including the case of flat patches), the chosen Dirichlet or Neumann condition fixes the far-field behavior of the solution.
In analogy to the spectral problem (3), the spectrum of the mixed Steklov-Neumann problem (23) is discrete, with a countable set of eigenvalues, which can be enumerated by the index $k$ in increasing order
$0=\mu_{0}^{N}<\mu_{1}^{N}\leq\mu_{2}^{N}\leq\ldots\nearrow+\infty.$ (24)
The restrictions of eigenfunctions onto the patch, $\Psi_{k}^{N}|_{\Gamma}$, form a complete orthonormal basis of $L^{2}(\Gamma)$. Note that the principal eigenvalue $\mu_{0}^{N}=0$ is associated to the constant eigenfunction $\Psi_{0}^{N}=1/\sqrt{|\Gamma|}$, and all other eigenfunctions are orthogonal to it.
The following spectral expansion was derived in [68]:
$\frac{1}{C(\mu)}=\frac{1}{C(\infty)}+\frac{2\pi}{\mu|\Gamma|}+2\pi\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}+\mu}\,.$ (25)
In this subsection, we discuss the complementary insights onto the reactive capacitance that can be gained from this expansion.
II.4 Sigmoidal approximation and several bounds
Neglecting the positive sum in Eq. (25), one gets immediately the lower bound:
$\frac{1}{C(\mu)}\geq\frac{1}{C(\infty)}+\frac{2\pi}{\mu|\Gamma|}=\frac{1}{C^{\rm app}(\mu)}\,.$ (26)
Rewriting the right-hand side of this bound as
$C^{\rm app}(\mu)=\frac{\mu\,C(\infty)}{\mu+2\pi C(\infty)/|\Gamma|}\,$, (27)
we retrieve the sigmoidal approximation for the reactive capacitance. This relation was proposed in [68] as an empirical approximation that correctly reproduces both small-$\mu$ and large-$\mu$ limits of $C(\mu)$. In turn, it follows here directly from the spectral expansion (25) by neglecting its last term, thus allowing one to control its accuracy (see below). Note that a similar approximation appeared in [90] in the analysis of the principal eigenvalue of the Laplace operator with Robin boundary condition, as well as in [37, 65]. In the special case of a circular patch, the approximation (27) was shown to be accurate over the entire range of $\mu$ within the maximal relative error of $4\%$ [68]. One of the aims of the present work consists in extending this result to patches of arbitrary shape. For this purpose, we need to compute numerically the Steklov eigenvalues $\mu_{k}$ and the spectral weights $F_{k}$ (see Sec. III).
To obtain an upper bound on $1/C(\mu)$, one can rewrite Eq. (25) as
$\frac{1}{C(\mu)}$ $=\frac{1}{C(\infty)}+\frac{2\pi}{\mu|\Gamma|}+2\pi\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}}$ (28)
$-2\pi\mu\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}(\mu_{k}^{N}+\mu)}\,.$
Taking the limit $\mu\to 0$ and using the Taylor expansion (17) with two terms, $C(\mu)\approx c_{1}\mu-c_{2}\mu^{2}+O(\mu^{3})$, one can compute the third term in the right-hand side of Eq. (28) as
$\frac{1}{C(\infty)}+2\pi\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}}$ $=\lim\limits_{\mu\to 0}\biggl(\frac{1}{C(\mu)}-\frac{2\pi}{\mu|\Gamma|}\biggr)$
$=\frac{c_{2}}{c_{1}^{2}}=2\pi\mathcal{A}_{\Gamma},$ (29)
where we used $c_{1}=|\Gamma|/(2\pi)$ and expressed $c_{2}$ in terms of the constant $\mathcal{A}_{\Gamma}$ defined in Eq. (20). As a consequence, Eq. (28) reads
$\frac{1}{C(\mu)}=2\pi\mathcal{A}_{\Gamma}+\frac{2\pi}{\mu|\Gamma|}-2\pi\mu\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}(\mu_{k}^{N}+\mu)}\,$, (30)
that yields the upper bound
$\frac{1}{C(\mu)}\leq 2\pi\mathcal{A}_{\Gamma}+\frac{2\pi}{\mu|\Gamma|}\,.$ (31)
The two bounds can be written together as
$\frac{1}{C(\infty)}\leq\frac{1}{C(\mu)}-\frac{2\pi}{\mu|\Gamma|}\leq 2\pi\mathcal{A}_{\Gamma}.$ (32)
We also note that the sum evaluated in Eq. (29) gives the upper bound on the absolute error of the sigmoidal approximation (27):
$\frac{1}{C(\mu)}-\frac{1}{C^{\rm app}(\mu)}$ $=2\pi\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}+\mu}$ (33)
$\leq 2\pi\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}}=2\pi\mathcal{A}_{\Gamma}-\frac{1}{C(\infty)}\,.$
Dividing this relation by $C^{\rm app}(\mu)$, we get the maximal bound on the relative error:
$\frac{C^{\rm app}(\mu)-C(\mu)}{C(\mu)}\leq\frac{{\mathcal{E}}_{\rm max}}{1+2\pi C(\infty)/(\mu|\Gamma|)}\,$, (34)
with
${\mathcal{E}}_{\rm max}=2\pi\mathcal{A}_{\Gamma}C(\infty)-1.$ (35)
One sees that the upper bound monotonously increases with $\mu$ and reaches the maximum ${\mathcal{E}}_{\rm max}$ at $\mu=\infty$. However, the actual relative error decreases in the limit $\mu\to\infty$ because $C(\mu)$ approaches to its limit $C(\infty)$, and the sigmoidal approximation correctly captures this behavior. As a consequence, the value ${\mathcal{E}}_{\rm max}$ is expected to be a conservative over-estimation. Using the spectral expansions (17, 21), one can represent the maximal bound as
${\mathcal{E}}_{\rm max}=\biggl(\sum\limits_{k=0}^{\infty}\mu_{k}F_{k}\biggr)\biggl(\sum\limits_{k=0}^{\infty}\frac{F_{k}}{\mu_{k}}\biggr)-1.$ (36)
According to Eq. (15), one sees that if $F_{0}$ is close to $1$, the dominant contributions to two sums are $\mu_{0}F_{0}$ and $F_{0}/\mu_{0}$ respectively, and the maximal bound is close to 0.
The bounds (32) can be further improved. For instance, rewriting Eq. (30) as
$\frac{1}{C(\mu)}$ $=2\pi\mathcal{A}_{\Gamma}+\frac{2\pi}{\mu|\Gamma|}-2\pi\mu\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{[\mu_{k}^{N}]^{2}}$
$+2\pi\mu^{2}\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{[\mu_{k}^{N}]^{2}(\mu_{k}^{N}+\mu)}\,$, (37)
one can neglect the last sum to obtain the improved lower bound on $1/C(\mu)$:
$\frac{1}{C(\mu)}\geq 2\pi\mathcal{A}_{\Gamma}+\frac{2\pi}{\mu|\Gamma|}-\mu B,$ (38)
where the constant
$B=2\pi\sum\limits_{k=1}^{\infty}\frac{[\Psi_{k}^{N}(\infty)]^{2}}{[\mu_{k}^{N}]^{2}}=\frac{c_{2}^{2}-c_{1}c_{3}}{c_{1}^{3}}$ (39)
was expressed in terms of $c_{i}$ from Eq. (18) via the Taylor expansion (17). Repeating this trick, one can get more and more accurate bounds that include, however, higher-order Taylor coefficients $c_{n}$.
II.5 Probabilistic interpretations
From a physical point of view, the reactive capacitance is proportional to the flux of particles diffusing from infinity, as stated earlier. In this subsection, we provide several probabilistic interpretations of $C(\mu)$ that will allow us to establish its additional properties.
Let $\bm{W}_{t}$ be the standard Brownian motion in ${\mathbb{R}}^{3}$ with diffusion coefficient $D$, that starts from a fixed point $\bm{y}$ at time 0. The reflected Brownian motion $\bm{X}_{t}$ in the upper half-space ${\mathbb{R}}^{3}_{+}$ can simply be obtained by reflecting any random path $\bm{W}_{t}$ with respect to the horizontal plane $y_{3}=0$; in other words, $\bm{X}_{t}=(W_{t}^{1},W_{t}^{2},|W_{t}^{3}|)$, where $W_{t}^{i}$ are three independent components of $\bm{W}_{t}$. To incorporate partial reactivity of the patch $\Gamma$, we follow the encounter-based approach [64]. For this purpose, we introduce the boundary local time $\ell_{t}$ on $\Gamma$ as the limit of the rescaled residence time $\ell_{t}^{(\sigma)}$ of $\bm{X}_{t}$, up to time $t$, in a thin boundary layer of thickness $\sigma$ near $\Gamma$:
$\ell_{t}=\lim\limits_{\sigma\to 0}\ell_{t}^{(\sigma)},\quad\ell_{t}^{(\sigma)}=\frac{D}{\sigma}\int\limits_{0}^{t}\Theta(\sigma-|\bm{X}_{t^{\prime}}-\Gamma|)dt^{\prime},$ (40)
where $\Theta(z)$ is the Heaviside step function: $\Theta(z)=1$ for $z>0$, and 0 otherwise. We note that both $\ell_{t}$ and $\ell_{t}^{(\sigma)}$ have units of length, despite their name. The boundary local time $\ell_{t}$ can also be interpreted as the rescaled number of arrivals of $\bm{X}_{t}$ onto $\Gamma$ up to time $t$. By construction, $\ell_{t}$ is a nondecreasing stochastic process, with $\ell_{0}=0$. Since the reflected Brownian motion in ${\mathbb{R}}^{3}_{+}$ is transient, its ability to revisit the patch is limited so that the boundary local time $\ell_{t}$ has a finite (random) limit $\ell_{\infty}$ as $t\to\infty$. Let us denote its probability density by $\rho_{\infty}(\ell|\bm{y})$.
At each arrival onto the patch $\Gamma$, the diffusing particle modeled by reflected Brownian motion may react and thus disappear. For a constant reactivity parameter $\mu$, the reaction event is triggered when the number of arrivals (i.e., failed reaction attempts), represented by $\ell_{t}$, exceeds a random threshold $\hat{\ell}$ that obeys the exponential distribution: ${\mathbb{P}}\{\hat{\ell}>\ell\}=e^{-\mu\ell}$ (see [64] for details). In this way, the first-reaction time (FRT) on the patch $\Gamma$ is defined as the first-crossing time of the threshold $\hat{\ell}$ by $\ell_{t}$: $\tau=\inf\{t>0~:~\ell_{t}<\hat{\ell}\}$. Since the reflected Brownian motion in ${\mathbb{R}}^{3}_{+}$ is transient, there is a finite probability of escaping to infinity without reacting on the patch; in this case, the FRT is infinite: $\tau=+\infty$. In turn, the probability of the reaction event, ${\mathbb{P}}_{\bm{y}}\{\tau<+\infty\}$, is known to satisfy Eq. (1) and is thus given by $w(\bm{y};\mu)$. We conclude that the auxiliary function $w(\bm{y};\mu)$ admits a simple interpretation as the probability of reaction on the patch $\Gamma$ with the reactivity parameter $\mu$, when the starting point is $\bm{y}$. In turn, $1-w(\bm{y};\mu)$ is the escape probability without reaction. We also note that the reaction time $\tau$ is finite if and only if the limit $\ell_{\infty}$ is above the threshold $\hat{\ell}$, i.e., the reaction event $\{\tau<+\infty\}$ is identical to $\{\ell_{\infty}>\hat{\ell}\}$. As a consequence, we get
$w(\bm{y};\mu)$ $={\mathbb{P}}_{\bm{y}}\{\tau<+\infty\}={\mathbb{P}}_{\bm{y}}\{\ell_{\infty}>\hat{\ell}\}$ (41)
$=\int\limits_{0}^{\infty}d\mu\,\underbrace{\mu e^{-\mu\ell}}_{\textrm{PDE of}~\hat{\ell}}~\cdot~\underbrace{{\mathbb{P}}_{\bm{y}}\{\ell_{\infty}>\ell\}}_{\int\nolimits_{\ell}^{\infty}\rho_{\infty}(\ell^{\prime}|\bm{y})d\ell^{\prime}}$
$=1-\int\limits_{0}^{\infty}d\mu\,e^{-\mu\ell}\,\rho_{\infty}(\ell|\bm{y})=1-{\mathbb{E}}_{\bm{y}}\{e^{-\mu\ell_{\infty}}\},$
where we integrated by parts. We conclude that $1-w(\bm{y};\mu)$ is the moment-generating function of the random variable $\ell_{\infty}$.
In the last step, we consider the situation when the starting position of the particle is not fixed but randomly chosen on the patch with the uniform density. The associated escape probability reads
$P_{\circ}(\mu)=\int\limits_{\Gamma}(1-w(\bm{y};\mu))\frac{d\bm{y}}{|\Gamma|}=\frac{2\pi C(\mu)}{\mu|\Gamma|}\,$, (42)
where we used the boundary condition (1b), the divergence theorem, and the asymptotic behavior (1d) at infinity (the subscript $\circ$ refers to the uniform distribution of the starting point on $\Gamma$). Inverting this relation, one can express the reactive capacitance $C(\mu)$ in terms of the escape probability on the patch. In turn, using Eq. (41), we can relate $C(\mu)$ to the moment-generating function:
${\mathbb{E}}_{\circ}\{e^{-\mu\ell_{\infty}}\}=\frac{2\pi C(\mu)}{\mu|\Gamma|}\,$, (43)
where
${\mathbb{E}}_{\circ}\{e^{-\mu\ell_{\infty}}\}=\int\limits_{\Gamma}{\mathbb{E}}_{\bm{y}}\{e^{-\mu\ell_{\infty}}\}\,\frac{d\bm{y}}{|\Gamma|}$ (44)
corresponds to the uniform starting point on the patch. Comparing this expression with the Taylor expansion (17), we get an interesting probabilistic interpretation of the Taylor coefficients $c_{n}$ as the moments of the boundary local time $\ell_{\infty}$:
${\mathbb{E}}_{\circ}\{\ell_{\infty}^{n}\}=\frac{2\pi}{|\Gamma|}c_{n+1}\quad(n=0,1,\ldots).$ (45)
For instance, we retrieve $c_{1}=|\Gamma|/(2\pi)$ at $n=0$, whereas $c_{2}$ is proportional to the mean boundary local time. Moreover, substituting the spectral expansion (14) into Eq. (43) and performing the inverse Laplace transform with respect to $\mu$, one gets the probability density of $\ell_{\infty}$:
$\rho_{\infty}(\ell|\circ)=\sum\limits_{k=0}^{\infty}\mu_{k}F_{k}e^{-\mu_{k}\ell}$ (46)
(see [64, 91] for further discussions on such spectral representations).
An alternative probabilistic interpretation of $C(\mu)$ can be deduced by multiplying Eq. (1a) by $w(\bm{y};\infty)$, Eq. (1a) with $\mu=\infty$ by $w(\bm{y};\mu)$, subtracting and integrating these equations over $\bm{y}\in{\mathbb{R}}^{3}_{+}$, and using the Green’s formula, boundary conditions and the behavior at infinity:
$\int\limits_{\Gamma}w(\bm{y};\mu)\partial_{n}w(\bm{y};\infty)d\bm{y}=\int\limits_{\Gamma}\partial_{n}w(\bm{y};\mu)d\bm{y}=2\pi C(\mu).$ (47)
Dividing this expression by $2\pi C(\infty)$, we get
$P_{h}(\mu)=\int\limits_{\Gamma}w(\bm{y};\mu)h(\bm{y})d\bm{y}=\frac{C(\mu)}{C(\infty)}\,$, (48)
where
$h(\bm{y})=\frac{1}{2\pi C(\infty)}\partial_{n}w(\bm{y};\infty)|_{\Gamma}$ (49)
can be interpreted as the (conditional) harmonic measure density for a particle arriving from infinity, i.e., $h(\bm{y})d\bm{y}$ is the probability of the first arrival onto the patch $\Gamma$ in a $d\bm{y}$ vicinity of the point $\bm{y}\in\Gamma$. In other words, $h(\bm{y})$ is the normalized flux density of particles from infinity. In this light, the integral in Eq. (48) can be interpreted as the probability of reaction for the process that has managed to arrive from infinity onto the patch.
II.6 Domain monotonicity
The electrostatic capacitance $C(\infty)$ is known to satisfy the domain monotonicity property (see [70, 78] for details):
$\textrm{If}~\Gamma_{1}\subset\Gamma_{2}\quad\Rightarrow\quad C_{1}(\infty)\leq C_{2}(\infty),$ (50)
i.e., the larger patch has the larger trapping ability. This property is an immediate consequence of the classical variational formulation of the electrostatic capacitance, which can be defined for compact sets (including the case of flat patches) as
$C(\infty)=\frac{1}{4\pi}\inf\biggl\{\int\limits_{{\mathbb{R}}^{3}}|\nabla u|^{2}d\bm{y}~:~u\in C_{0}^{\infty}({\mathbb{R}}^{3}),~u|_{\Gamma}\geq 1\biggr\},$ (51)
where $C^{\infty}_{0}({\mathbb{R}}^{3})$ is the space of smooth functions in ${\mathbb{R}}^{3}$ with a compact support, and the factor $4\pi$ is included to ensure our convention that the electrostatic capacitance of a ball of radius $R$ is $R$; a similar definition can be written for the upper half-space. Note that the functional space $C_{0}^{\infty}$ can be replaced by the Sobolev space $H^{1}({\mathbb{R}}^{3})$ with appropriate restrictions.
It is intuitively appealing that the same statement should hold for partially reactive patches, i.e., a larger patch captures diffusing particles more efficiently. In this subsection, we establish this domain monotonicity property and some other monotonicity relations.
First monotonicity relation
For this purpose, we first derive another representation of $C(\mu)$ by setting $\bar{w}(\bm{y};\mu)=1-w(\bm{y};\mu)$, which satisfies
$\Delta\bar{w}$ $=0\quad(\bm{y}\in{\mathbb{R}}^{3}_{+}),$ (52a)
$\partial_{n}\bar{w}+\mu\bar{w}$ $=0\quad(\bm{y}\in\Gamma),$ (52b)
$\partial_{n}\bar{w}$ $=0\quad(\bm{y}\in\partial{\mathbb{R}}^{3}_{+}\backslash\overline{\Gamma}),$ (52c)
$\bar{w}$ $\sim 1-\frac{C(\mu)}{|\bm{y}|}+\mathcal{O}(|\bm{y}|^{-2})\quad(|\bm{y}|\to\infty).$ (52d)
Multiplying this equation by $\bar{w}$ and integrating over ${\mathbb{R}}^{3}_{+}$, we get
$0$ $=\int\limits_{{\mathbb{R}}^{3}_{+}}\bar{w}\,\Delta\bar{w}\,d\bm{y}=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\bar{w}|^{2}\,d\bm{y}+\int\limits_{\Gamma}\bar{w}\,\partial_{n}\bar{w}\,d\bm{y}$
$=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\bar{w}|^{2}\,d\bm{y}-\mu\int\limits_{\Gamma}\bar{w}^{2}\,d\bm{y}+2\pi C(\mu),$
where we used the Green’s formula to integrate by parts, the Robin boundary condition (52b), and the behavior (52d) at infinity. As a consequence, we have
$C(\mu)=\frac{1}{2\pi}\biggl\{\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla w|^{2}\,d\bm{y}+\mu\int\limits_{\Gamma}(1-w)^{2}\,d\bm{y}\biggr\},$ (53)
where we returned to the original function $w(\bm{y};\mu)$.
Following the standard arguments, one can recast this relation in a variational form:
$C(\mu)=\frac{1}{2\pi}\inf\limits_{u\in C^{\infty}_{0}({\mathbb{R}}^{3}_{+})}\biggl\{\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla u|^{2}d\bm{y}+\mu\int\limits_{\Gamma}(1-u)^{2}\,d\bm{y}\biggr\}.$ (54)
The variational formulation immediately implies the domain monotonicity property for the reactive capacitance:
$\textrm{If}~\Gamma_{1}\subset\Gamma_{2}\quad\Rightarrow\quad C_{1}(\mu)\leq C_{2}(\mu)~~~\forall~\mu\geq 0,$ (55)
where $C_{1}(\mu)$ and $C_{2}(\mu)$ refer to the reactive capacitances of $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Indeed, as one searches the infimum over the same functional space, any reduction of the patch diminishes the second term in Eq. (54) and thus the associated $C_{i}(\mu)$. This is an extension of Eq. (50) to a finite reactivity $\mu>0$.
Second monotonicity relation
In the same vein, one can multiply Eq. (1a) by $w(\bm{y};\mu)$ and then integrating over ${\mathbb{R}}^{3}_{+}$ to get:
$0$ $=\int\limits_{{\mathbb{R}}^{3}_{+}}w\,\Delta w\,d\bm{y}=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla w|^{2}\,d\bm{y}+\int\limits_{\Gamma}w\,\partial_{n}w\,d\bm{y}$
$=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla w|^{2}\,d\bm{y}+\mu(\mu|\Gamma|-2\pi C(\mu))-\mu\int\limits_{\Gamma}w^{2}\,d\bm{y},$
where we used again the Green’s formula, the boundary condition (1b), and the behavior (1d) at infinity such that
$\mu\int\limits_{\Gamma}(1-w)d\bm{y}=\int\limits_{\Gamma}\partial_{n}w\,d\bm{y}=2\pi C(\mu).$ (56)
We have thus
$\delta C(\mu)=\frac{1}{2\pi\mu}\biggl\{\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla w|^{2}\,d\bm{y}+\mu\int\limits_{\Gamma}w^{2}\,d\bm{y}\biggr\},$ (57)
with a shortcut notation
$\delta C(\mu)=\frac{\mu|\Gamma|}{2\pi}-C(\mu).$ (58)
The variational form reads
$\delta C(\mu)=\frac{1}{2\pi\mu}\inf\limits_{u\in C^{\infty}_{0}({\mathbb{R}}^{3}_{+})}\biggl\{\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla u|^{2}d\bm{y}+\mu\int\limits_{\Gamma}u^{2}\,d\bm{y}\biggr\}$ (59)
and thus yields the second domain monotonicity property:
$\textrm{If}~\Gamma_{1}\subset\Gamma_{2}\quad\Rightarrow\quad\delta C_{1}(\mu)\leq\delta C_{2}(\mu)~~~\forall~\mu\geq 0.$ (60)
This property could alternatively be obtained from the probabilistic interpretation. For any fixed $\bm{y}$ and $\mu$, the boundary local time on the larger patch $\Gamma_{2}$ is greater than or equal to the boundary local time on the smaller patch, so that the first-reaction time on $\Gamma_{2}$ is smaller and thus the particle has more chances to react with $\Gamma_{2}$ before escaping to infinity. As a consequence, one has
$\textrm{If}~\Gamma_{1}\subset\Gamma_{2}\quad\Rightarrow\quad 0\leq w_{1}(\bm{y};\mu)\leq w_{2}(\bm{y};\mu)~~~\left(\mu\geq 0,\atop\bm{y}\in{\mathbb{R}}^{3}_{+}\right).$ (61)
The integral of this inequality over $\Gamma_{1}$ yields then
$\int\limits_{\Gamma_{1}}w_{1}(\bm{y};\mu)d\bm{y}\leq\int\limits_{\Gamma_{1}}w_{2}(\bm{y};\mu)d\bm{y}\leq\int\limits_{\Gamma_{2}}w_{2}(\bm{y};\mu)d\bm{y},$
from which Eq. (56) implies
$C_{2}(\mu)\leq C_{1}(\mu)+\frac{|\Gamma_{2}|-|\Gamma_{1}|}{2\pi}\mu\,.$ (62)
Third monotonicity relation
Next, we establish another monotonicity relation. For this purpose, let us consider the auxiliary function
$\tilde{w}(\bm{y};\mu)=\frac{w(\bm{y};\mu)}{C(\mu)}-\frac{w(\bm{y};\infty)}{C(\infty)}\,$, (63)
which satisfies, by construction,
$\Delta\tilde{w}$ $=0\quad(\bm{y}\in{\mathbb{R}}^{3}_{+}),$ (64a)
$\partial_{n}\tilde{w}+\mu\tilde{w}$ $=\frac{\mu}{C(\mu)}-\frac{\mu}{C(\infty)}-2\pi h(\bm{y})\quad(\bm{y}\in\Gamma),$ (64b)
$\partial_{n}\tilde{w}$ $=0\quad(\bm{y}\in\partial{\mathbb{R}}^{3}_{+}\backslash\overline{\Gamma}),$ (64c)
$\tilde{w}$ $\sim\mathcal{O}(|\bm{y}|^{-2})\quad(|\bm{y}|\to\infty),$ (64d)
where $h(\bm{y})$ was defined in Eq. (49). Since $\tilde{w}$ decays faster than $1/|\bm{y}|$, the divergence theorem implies
$\int\limits_{\Gamma}\partial_{n}\tilde{w}=0,$ (65)
from which the boundary condition (64b) yields
$\int\limits_{\Gamma}\tilde{w}(\bm{y};\mu)d\bm{y}=\frac{|\Gamma|}{C(\mu)}-\frac{|\Gamma|}{C(\infty)}-\frac{2\pi}{\mu}\,$, (66)
where we used that $\int\nolimits_{\Gamma}h(\bm{y})\,d\bm{y}=1$. Multiplying Eq. (64a) by $\tilde{w}$ and integrating over ${\mathbb{R}}^{3}_{+}$, we get
$0$ $=\int\limits_{{\mathbb{R}}^{3}_{+}}\tilde{w}\,\Delta\tilde{w}=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\tilde{w}|^{2}+\int\limits_{\Gamma}\tilde{w}\,\partial_{n}\tilde{w}$.
Substituting $\partial_{n}\tilde{w}$ from the boundary condition (64b), we have
$0$ $=-\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\tilde{w}|^{2}-\mu\int\limits_{\Gamma_{i}}\tilde{w}^{2}+\biggl(\frac{\mu}{C(\mu)}-\frac{\mu}{C(\infty)}\biggr)\int\limits_{\Gamma}\tilde{w}$
$-2\pi\int\limits_{\Gamma}\tilde{w}\,h\,.$
To evaluate the last term, we multiply Eq. (64a) by $w(\bm{y};\infty)$, multiply Eq. (1a) with $\mu=\infty$ by $\tilde{w}(\bm{y};\mu)$, subtract them, integrate over ${\mathbb{R}}^{3}_{+}$, use the Green’s formula, boundary conditions and the decay at infinity to get
$\int\limits_{\Gamma}\tilde{w}\,h=\int\limits_{\Gamma}\underbrace{w(\bm{y};\infty)}_{=1}\partial_{n}\tilde{w}=0,$ (67)
where we used Eq. (65). Finally, substituting Eq. (66), we obtain
$\mu|\Gamma|\biggl(\frac{1}{C(\mu)}-\frac{1}{C(\infty)}\biggr)\biggl(\frac{1}{C(\mu)}-\frac{1}{C(\infty)}-\frac{2\pi}{\mu|\Gamma|}\biggr)$
$=\int\limits_{{\mathbb{R}}^{3}_{+}}|\nabla\tilde{w}|^{2}+\mu\int\limits_{\Gamma}\tilde{w}^{2}.$ (68)
Using again the variational form, we obtain the following inequality
$|\Gamma_{1}|\biggl(\frac{1}{C_{1}(\mu)}-\frac{1}{C_{1}(\infty)}\biggr)\biggl(\frac{1}{C_{1}(\mu)}-\frac{1}{C_{1}(\infty)}-\frac{2\pi}{\mu|\Gamma_{1}|}\biggr)$
$\leq|\Gamma_{2}|\biggl(\frac{1}{C_{2}(\mu)}-\frac{1}{C_{2}(\infty)}\biggr)\biggl(\frac{1}{C_{2}(\mu)}-\frac{1}{C_{2}(\infty)}-\frac{2\pi}{\mu|\Gamma_{2}|}\biggr)$ (69)
if $\Gamma_{1}\subset\Gamma_{2}$.
Fourth monotonicity relation
From the last inequality, let us deduce a weaker yet more appealing monotonicity property. Denoting $\eta_{i}=|\Gamma_{i}|(1/C_{i}(\mu)-1/C_{i}(\infty))$, we rewrite Eq. (69) as
$\frac{\eta_{1}(\eta_{1}-2\pi/\mu)}{|\Gamma_{1}|}\leq\frac{\eta_{2}(\eta_{2}-2\pi/\mu)}{|\Gamma_{2}|}$ (70)
or, equivalently,
$\frac{(\eta_{1}-\pi/\mu)^{2}-\pi^{2}/\mu^{2}}{|\Gamma_{1}|}\leq\frac{(\eta_{2}-\pi/\mu)-\pi^{2}/\mu^{2}}{|\Gamma_{2}|}\,.$ (71)
Since $|\Gamma_{1}|\leq|\Gamma_{2}|$, multiplication of two inequalities yields, after few simplifications,
$|\eta_{1}-\pi/\mu|\leq|\eta_{2}-\pi/\mu|\,.$ (72)
Since $\eta_{i}\geq 2\pi/\mu$ due to Eq. (25), we conclude that $\eta_{1}\leq\eta_{2}$, i.e.,
$|\Gamma_{1}|\biggl(\frac{1}{C_{1}(\mu)}-\frac{1}{C_{1}(\infty)}\biggr)\leq|\Gamma_{2}|\biggl(\frac{1}{C_{2}(\mu)}-\frac{1}{C_{2}(\infty)}\biggr).$ (73)
This inequality allows one to control the large-reactivity behavior of the reactive capacitance.
II.7 Large-reactivity limit
When the reactivity parameter $\mu$ goes to infinity, the reactive capacitance $C(\mu)$ approaches its limit $C(\infty)$. However, the behavior of the reactive capacitance is not analytic at infinity. In fact, according to Eq. (2), the reactive capacitance is the integral of the flux density, which may diverge at the boundary of the patch in the Dirichlet setting. This is well-known for an elliptic patch with semiaxes $a\leq b$, for which the flux density is given by (see, e.g., [94])
$\partial_{n}w(\bm{y};\infty)|_{\Gamma}=\frac{\bigl(1-y_{1}^{2}/a^{2}-y_{2}^{2}/b^{2}\bigr)^{-1/2}}{aK(\sqrt{1-a^{2}/b^{2}})}\,$, (74)
where $K(z)$ is the complete elliptic integral of the first kind:
$K(z)=\int\limits_{0}^{\pi/2}\frac{d\theta}{\sqrt{1-z^{2}\sin^{2}\theta}}\,.$ (75)
The effect of this singular behavior onto the mean first-reaction time on a circular patch was analyzed in [44] by using a Wiener-Hopf integral equation. Relying on this analysis, the large-$\mu$ asymptotic behavior of the reactive capacitance for the unit disk was deduced in [68]:
$C_{\circ}(\mu)\approx C_{\circ}(\infty)-\frac{2(\ln(2\mu)+\gamma+1)}{\pi^{2}\mu}\quad(\mu\to\infty),$ (76)
where $\gamma\approx 0.5772\ldots$ is the Euler constant, $C_{\circ}(\infty)=2/\pi$, and subscript $\circ$ refers to the unit disk. The presence of the logarithm makes this expression non-analytic. Does this behavior hold for noncircular patches?
While a systematic analysis of this problem for arbitrary patches remains an open problem, we argue here that the non-analytic large-$\mu$ behavior is a generic feature of flat patches. For this purpose, we first rewrite Eq. (76) as
$\frac{1}{C_{\circ}(\mu)}-\frac{1}{C_{\circ}(\infty)}\approx\frac{\ln(2\mu)+\gamma+1}{2\mu}\quad(\mu\to\infty).$ (77)
This relation can now be used as a bound according to the monotonicity relation (73). In fact, for any given patch $\Gamma$, one can inscribe a small disk $\Gamma_{0}$ of radius $r_{0}$. As $\Gamma_{0}\subset\Gamma$, one has
$\frac{\pi r_{0}}{|\Gamma|}\biggl(\frac{1}{C_{\circ}(r_{0}\mu)}-\frac{1}{C_{\circ}(\infty)}\biggr)\leq\frac{1}{C(\mu)}-\frac{1}{C(\infty)}\,$, (78)
where we used the rescaling relation (22). Even though the asymptotic relation (77) is not a bound, one can add an additional prefactor to turn it into a bound for large $\mu$, i.e.,
$\beta\frac{\ln(2r_{0}\mu)+\gamma+1}{2\mu}\leq\frac{1}{C(\mu)}-\frac{1}{C(\infty)}\quad(\mu\to\infty),$ (79)
with some constant $\beta>0$. We conclude that the difference in the right-hand side cannot vanish faster than $\ln(\mu)/\mu$, i.e., the logarithmic singularity is also expected at infinity, as for the circular patch. A further work is needed to characterize this behavior more accurately.
The non-analytic behavior of $C(\mu)$ is tightly related to the divergence of the sum of coefficients $V_{k}^{N}(\infty)]^{2}$. In fact, the above asymptotic behavior, together with the spectral expansion (25), implies that
$\mu\biggl(\frac{1}{C(\mu)}-\frac{1}{C(\infty)}\biggr)=2\pi\sum\limits_{k=0}^{\infty}\frac{\mu[\Psi_{k}^{N}(\infty)]^{2}}{\mu_{k}^{N}+\mu}\propto\ln(\mu)\to\infty$ (80)
as $\mu\to\infty$. This divergence suggests that their partial sum also diverges logarithmically:
$\sum\limits_{k=1}^{K}[\Psi_{k}^{N}(\infty)]^{2}\propto\ln(K)\qquad(K\to\infty).$ (81)
It is instructive to give an alternative insight. For this purpose, let us multiply Eq. (23a) by $w(\bm{y};\infty)$, Eq. (1a) by $\Psi_{k}^{N}(\bm{y})$, subtract them, integrate over ${\mathbb{R}}^{3}_{+}$ and use the boundary conditions and the asymptotic behavior at infinity to get
$\Psi_{k}^{N}(\infty)=\int\limits_{\Gamma}h(\bm{y})\,\Psi_{k}^{N}(\bm{y})d\bm{y},$ (82)
where $h(\bm{y})$ is defined by Eq. (49). Since $\{\Psi_{k}^{N}\}$ form a complete basis of $L^{2}(\Gamma)$, one can formally write
$\sum\limits_{k=0}^{\infty}[\Psi_{k}^{N}(\infty)]^{2}$ $=\int\limits_{\Gamma}d\bm{y}_{1}\,h(\bm{y}_{1})\int\limits_{\Gamma}d\bm{y}_{2}\,h(\bm{y}_{2})$
$\times\underbrace{\sum\limits_{k=0}^{\infty}\Psi_{k}^{N}(\bm{y}_{1})\Psi_{k}^{N}(\bm{y}_{2})}_{=\delta(\bm{y}_{1}-\bm{y}_{2})}=\int\limits_{\Gamma}d\bm{y}\,h^{2}(\bm{y})=\infty.$
In other words, the squared harmonic measure density diverges logarithmically due to the edge singularity. This is exemplified by Eq. (74) that provides an explicit form of $h(\bm{y})$ for elliptic patches. In summary, even though the coefficients $\Psi_{k}^{N}(\infty)]^{2}$ determine the relative contributions of the Steklov eigenfunctions $\Psi_{k}^{N}$ to the reactive capacitance via Eq. (25), in analogy to the coefficients $F_{k}$ in Eq. (14), the latter are summed to $1$ that facilitates their interpretation as relative weights of the Steklov eigenfunctions $\Psi_{k}$.
II.8 Bounds for the principal eigenvalue
While there were numerous studies on isoperimetric inequalities for Steklov eigenvalues in the conventional setting of interior domains (see [74, 75, 76] and references therein), much less is known about the exterior Steklov problem [92, 73]. Moreover, the assumptions on the regularity of the boundary that were used in former works, do not necessarily hold for flat patches. For instance, a circular patch can be seen as the limit of thinning oblate spheroids, whose curvature diverges (the edge singularity). To fulfill this gap, we briefly mention some inequalities that follow from our spectral expansions, without pretending for rigorous proofs.
Since $\mu_{k}$ and $F_{k}$ are positive, one can use the spectral expansion (21) to recover the Payne’s upper bound for the principal eigenvalue $\mu_{0}$ [93]:
$\mu_{0}\leq\frac{2\pi C(\infty)}{|\Gamma|}\,.$ (83)
In a similar way, one can substitute $\mu_{0}<\mu_{k}$ into the spectral representations (17) of the Taylor coefficients $c_{n}$ to get
$\mu_{0}\leq\biggl(\frac{|\Gamma|}{2\pi c_{n+1}}\biggr)^{1/n}\,,\qquad n=1,2,\ldots.$ (84)
Moreover, keeping only the first term in the spectral expansion (17), one gets the lower bound:
$\mu_{0}\geq\biggl(\frac{F_{0}\,|\Gamma|}{2\pi c_{n+1}}\biggr)^{1/n}\,,\qquad n=1,2,\ldots,$ (85)
which is less explicit as it includes the spectral coefficient $F_{0}$. For instance, one has for $n=1$
$\frac{F_{0}}{\mathcal{A}_{\Gamma}|\Gamma|}\leq\mu_{0}\leq\frac{1}{\mathcal{A}_{\Gamma}|\Gamma|}\,$, (86)
with $\mathcal{A}_{\Gamma}$ being defined in Eq. (20). Alternatively, the inequality (85) can be seen as an upper bound on $F_{0}$:
$F_{0}\leq\frac{2\pi c_{n+1}\mu_{0}^{n}}{|\Gamma|}\,,\qquad n=1,2,\ldots,$ (87)
or, equivalently,
$\int\limits_{\Gamma}\Psi_{0}\leq\sqrt{2\pi c_{n+1}\mu_{0}^{n}}\,,\qquad n=1,2,\ldots$ (88)
To illustrate these bounds, let us consider the circular patch of unit radius, for which the first three Taylor coefficients were obtained in [68]:
$c_{1}=\frac{1}{2},\quad c_{2}=\frac{4}{3\pi},\quad c_{3}=\frac{4}{\pi^{2}}\int\limits_{0}^{1}xE^{2}(x)dx\approx 0.3651,$ (89)
where $E(x)$ is the complete elliptic integral of the second kind. Substituting the values of $C(\infty)=2/\pi$, $c_{2}$ and $c_{3}$ into Eqs. (83, 84), we get three upper bounds: $1.263$ (Payne’s bound), $1.178$ ($n=1$), and $1.170$ ($n=2$), whereas the numerical value of the principal eigenvalue is $1.159$. One sees that the bound becomes tighter as $n$ increases, as expected from the spectral form (21). Using the numerical value $F_{0}\approx 0.978$, we get the lower bounds $1.152$ and $1.157$, where the last value is very close to the numerical value of $\mu_{0}$. We recall that $c_{2}$ and $c_{3}$ for arbitrary patches are expressed via Eq. (18) in terms of the function $\omega_{\Gamma}(\bm{x})$ from Eq. (19). In Appendix B, we compute $\mathcal{A}_{\Gamma}$ for rectangular and elliptic patches; in the latter case, we also compare the upper bounds (83) and (86).
II.9 Relation to the patch shape
To gain further insights onto the relation between the reactive capacitance of the patch and its geometric shape, one needs to solve the exterior Steklov problem (3). In Appendix A, we develop an efficient numerical approach, which relies on the reformulation (11) of the original problem. This is a finite-element method, which employs linear basis functions on a triangular mesh on the patch. The key step to achieve good accuracy and rapidity of the method is a semi-analytical representation of the integral kernel $\mathcal{G}(\bm{y},\bm{y}^{\prime})$ on these basis functions. The accuracy of the method, which is controlled by the mesh size, was validated by considering the circular patch, for which an alternative very accurate approach based on oblate spheroidal coordinates is available [80].
Using this numerical tool, we aim at answering the following questions:
(i) How does anisotropy of the patch affect the Steklov eigenvalues and eigenfunctions?
(ii) Does the principal eigenfunction $\Psi_{0}$ always provide the dominant contribution to the reactive capacitance?
(iii) What is the accuracy of the empirical approximation (27) for different patches?
III Numerical results
III.1 Steklov eigenpairs
Since the Steklov eigenfunctions determine the reactive capacitance, it is instructive to first inspect their behavior for different patches. We start with the circular patch as a standard benchmark example, and then explore other patch shapes.
Circular patch
Figure 2: First 9 Steklov eigenfunctions $\Psi_{k}$ restricted to the circular patch of unit radius (the associated eigenvalues $\mu_{k}$ and weights $F_{k}$ are shown in the titles). These eigenfunctions were obtained by using a mesh with 688 triangles and 375 nodes.
Figure 2 presents the first 9 Steklov eigenfunctions $\Psi_{k}|_{\Gamma}$ on the circular patch of unit radius. The associated eigenvalues $\mu_{k}$ and weights $F_{k}$ are shown in the panels. As discussed in Appendix A.3 (see also [80]), the axial symmetry of the problem implies that the eigenfunctions can be written in polar coordinates $(r,\phi)$ as $e^{\pm im\phi}v_{m,n}(r)$, with positive integer $m$ and $n$. For any $m\neq 0$, a linear combination of two non-axially symmetric eigenfunctions $e^{im\phi}v_{m,n}(r)$ and $e^{-im\phi}v_{m,n}(r)$ is also an eigenfunction, implying that the associated eigenvalue is at least twice degenerate. This is confirmed by the shape of the shown eigenfunctions (which are chosen to be real-valued), e.g., $\Psi_{1}$ and $\Psi_{2}$ correspond to the same eigenvalue. Our numerical results suggest that the eigenvalues corresponding to axially symmetric eigenfunctions are simple, whereas the eigenvalues corresponding to non-axially symmetric eigenfunctions are twice degenerate. We stress that only axially symmetric eigenfunctions (here, with indices $k=0$ and $k=5$) do contribute to the reactive capacitance $C(\mu)$ because the projection of any non-axially symmetric eigenfunction onto a constant is zero.
Elliptic patches
Figure 3: First 9 Steklov eigenfunctions $\Psi_{k}$ restricted to the elliptic patch of semiaxes $0.5$ and 11 (the associated eigenvalues $\mu_{k}$ and weights $F_{k}$ are shown in the titles). The eigenfunctions were obtained by using a mesh with 756 triangles and 413 nodes.
The trapping capacity of a perfectly reactive elliptic patch was studied by Strieder [94]. In particular, the exact solution of Eq. (1) at $\mu=\infty$, that determines the concentration profile $w(\bm{y};\infty)$ in the upper half-space, was given in ellipsoidal coordinates (see also [95]), from which the flux density in Eq. (74) follows. Integrating this density over the patch, one retrieves the electrostatic capacitance [70]:
$C(\infty)=\frac{b}{K(\sqrt{1-a^{2}/b^{2}})}\,.$ (90)
Unfortunately, these analytical results do not admit explicit extensions to partially reactive elliptic patches. In turn, one can express the flux density and the reactive capacitance in terms of the Steklov eigenfunctions.
Figure 3 presents the first 9 eigenfunctions $\Psi_{k}|_{\Gamma}$ and related eigenvalues $\mu_{k}$ for an elliptic patch with semiaxes $1$ and $0.5$. In contrast to the circular patch, all shown eigenvalues are simple, i.e., anisotropy generally removes the degeneracy of eigenvalues. At the same time, the continuous dependence of the eigenvalues on the aspect ratio $a/b$ allows one to construct ellipses, for which some eigenvalues are degenerate (e.g., the shown eigenvalues $\mu_{2}$ and $\mu_{3}$ are close, and they can be made equal by varying $a/b$). The symmetry of the elliptic patch implies that some eigenfunctions are antisymmetric with respect to the horizontal or vertical axis, so that their projection onto a constant is zero. Such antisymmetric eigenfunctions do not contribute to the reactive capacitance. Among the shown eigenfunctions, only $\Psi_{0}$, $\Psi_{3}$, $\Psi_{7}$ and $\Psi_{8}$ do contribute to $C(\mu)$, with their relative weights: $F_{0}\approx 0.9741$, $F_{3}\approx 0.0138$, $F_{7}\approx 0.0074$ and $F_{8}\approx 10^{-5}$. One sees that, as for the circular patch, the first eigenfunction $\Psi_{0}$ provides the dominant contribution of $97\%$.
We complete this discussion by illustrating the difference between the Steklov problems (3) and (23). Figure 4 presents the first 9 eigenfunctions $\Psi_{k}^{N}|_{\Gamma}$ for the elliptic patch of semiaxes $1$ and $0.5$. Their comparison with panels of Fig. 3 indicates that, apart from the principal eigenfunctions $\Psi_{0}$ and $\Psi_{0}^{N}$, which are substantially different, the other eigenfunctions look quite similar, up to matching their indices. Moreover, antisymmetric eigenfunctions are actually identical, up to their signs, e.g. $\Psi_{1}^{N}=-\Psi_{1}$, $\Psi_{3}^{N}=-\Psi_{2}$, etc. This is not surprising (see Appendix A for details): as the kernels $\mathcal{G}(\bm{y},\bm{y}^{\prime})$ and $\mathcal{G}^{N}(\bm{y},\bm{y}^{\prime})$ differ by a function $\omega_{\Gamma}(\bm{y})/|\Gamma|+\omega_{\Gamma}(\bm{y}^{\prime})/|\Gamma|-\mathcal{A}_{\Gamma}$, which is symmetric on the elliptic patch, the antisymmetric eigenfunctions of these kernels should be identical.
Figure 4: First 9 Steklov eigenfunctions $\Psi_{k}^{N}$ restricted to the elliptic patch of semiaxes $0.5$ and $1$ for the Neumann version (the associated eigenvalues are shown in the titles). The eigenfunctions were obtained by using a mesh with 756 triangles and 413 nodes.
Rectangular patches
Figures 5 and 6 present similar results for a quadratic patch $(-1,1)\times(-1,1)$ and for an elongated rectangular patch $(-1,1)\times(-0.2,0.2)$. As in the case of a circular patch, symmetries of the square make some eigenvalues degenerate, e.g., $\mu_{1}=\mu_{2}$ and $\mu_{6}=\mu_{7}$. In turn, this degeneracy is generally removed for rectangular patches, even though continuous variations of $a/b$ can be used to obtained some degenerate eigenvalues.
Figure 5: First 9 Steklov eigenfunctions $\Psi_{k}$ restricted to the square patch $(-1,1)\times(-1,1)$ (the associated eigenvalues $\mu_{k}$ and weights $F_{k}$ are shown in the titles). The eigenfunctions were obtained by using a mesh with 674 triangles and 372 nodes.
Figure 6: First 9 Steklov eigenfunctions $\Psi_{k}$ restricted to the rectangular patch $(-1,1)\times(-0.2,0.2)$ (the associated eigenvalues $\mu_{k}$ and weights $F_{k}$ are shown in the titles). The eigenfunctions were obtained by using a mesh with 704 triangles and 400 nodes.
Disconnected patches
To give a broader view onto the spectral properties, we also discuss the case of disconnected patches. In sharp contrast to Laplacian eigenfunctions, for which the connectivity of the domain plays the crucial role, it is less relevant for the considered Steklov problem. For instance, splitting a patch into two disconnected subsets by a curve has no effect onto the spectrum. Moreover, since the kernel $\mathcal{G}(\bm{y},\bm{y}^{\prime})$ decays slowly with the distance between $\bm{y}$ and $\bm{y}^{\prime}$, even if two subsets are separated by a significant distance, they still compete with each other for capturing diffusing particles. This long-range interaction, known as diffusional screening [96, 97, 98, 99] or diffusive interaction [100], implies that the electrostatic capacitance is not additive, i.e., the capacitance of the union of two patches is always smaller than the sum of their individual capacitances. For perfect reactions, this competition for the case of two circular patches on the reflecting plane was characterized analytically in [101, 102], whereas Lindsay and Bernoff derived asymptotic formulas in the case of multiple small circular patches [29] (see also [36]). A further extension to partially reactive patches of arbitrary shape was given in [69]. It is therefore instructive to inspect the behavior of the Steklov eigenfunctions in the case of disconnected patches.
Figure 7 shows the first 15 Steklov eigenfunctions $\Psi_{k}|_{\Gamma}$ for two circular patches of radii $1$ and $0.5$, separated by distance $0.5$. Interestingly, the presence of the smaller patch has only a minor effect on the principal eigenvalue $\mu_{0}$ and the spectral weight $F_{0}$ of the associated eigenfunction. We also note that, despite the proximity of two patches, some eigenfunctions are localized on one patch (i.e., they almost vanish on the other). The refined structure of the Steklov eigenpairs gives access to the reactive capacitance and its dependence on the spatial arrangement of the disconnected patches. Further analysis of this problem presents an interesting perspective of this work.
Figure 7: First 15 Steklov eigenfunctions $\Psi_{k}$ restricted to the patch formed by two disks of radii $1$ and $0.5$ (the associated eigenvalues $\mu_{k}$ and weights $F_{k}$ are shown in the titles). The eigenfunctions were obtained by using a mesh with 964 triangles and 536 nodes.
III.2 Effect of patch anisotropy onto the eigenvalues
To investigate more quantitatively the role of patch anisotropy, we inspect three families of patches: (i) ellipses of semi-axes $a$ and $b$, (ii) rectangles $(-a,a)\times(-b,b)$, and (iii) rhombuses with diagonals $2a$ and $2b$. The patch surface areas are respectively $\pi ab$, $4ab$, and $2ab$. We fix $b=1$ and then change the anisotropy by varying $a$ from 0 to $1$. In the limit $a\to 0$, all three shapes approach the interval (1,1)(-1,1), which is inaccessible to three-dimensional Brownian motion and thus presents a singular limit.
We first inspect the asymptotic behavior of the eigenvalues $\mu_{k}$ for an elliptic patch as $a\to 0$. Since the elliptic patch shrinks to an interval, the eigenvalues are expected to diverge as $a\to 0$. The principal eigenvalue $\mu_{0}$ cannot diverge faster than the right-hand side of the Payne’s inequality (83). Substituting the asymptotic behavior
$K(z)\sim 2\ln(2)-\frac{1}{2}\ln(1-z^{2})\qquad(z\to 1)$ (91)
to Eq. (90), one gets
$C(\infty)\sim b/\ln(4b/a)\qquad(a\to 0).$ (92)
As a consequence, $\mu_{0}$ cannot grow faster than $\propto(a\ln(4b/a))^{-1}$. Figure 8 shows $1/(a\mu_{0})$ (circles) as a function of $\ln(b/a)$ to confirm this behavior. Indeed, one observes a linear dependence,
$\frac{1}{a\mu_{0}}\approx 0.59\ln(b/a)+0.71\qquad(a\ll b),$ (93)
with the coefficients obtained from a linear fit (shown by dashed line).
To our knowledge, there is no analytic expression for the capacitance of a rectangular patch. For this reason, we employ the asymptotic behavior (141) of the coefficient $\mathcal{A}_{\Gamma}$ for elongated rectangles (see Appendix B). According to the lower and upper bounds (86), the asymptotic behavior of $\mathcal{A}_{\Gamma}$ is expected to control that of the principal eigenvalue $\mu_{0}$. Substituting Eq. (141) into Eq. (86), we get approximately
$\mu_{0}\approx\frac{1}{|\Gamma|\mathcal{A}_{\Gamma}}\simeq\frac{\pi}{a(1+2\ln(2b/a))}\qquad(a\ll b).$ (94)
or, equivalently,
$\frac{1}{a\mu_{0}}\approx\frac{1+2\ln(2b/a)}{\pi}\approx 0.64\ln(b/a)+0.76\qquad(a\ll b).$ (95)
Figure 8 presents by squares the left-hand side of this relation, with $\mu_{0}$ obtained numerically. A linear fit of these values is shown by solid line and yields $1/(a\mu_{0})\approx 0.63\ln(b/a)+0.83$, in an excellent agreement with our prediction (95).
A similar asymptotic analysis could be performed for rhombic patches. However, we skip this analysis and just present a linear fit of $1/(a\mu_{0})$ versus $\ln(b/a)$, obtained from numerical values of $\mu_{0}$ and shown in Fig. 8 by dash-dotted line:
$\frac{1}{a\mu_{0}}\approx 0.52\ln(b/a)+0.45\qquad(a\ll b).$ (96)
In all three cases, the asymptotic behavior is similar but the coefficients are slightly different. We outline that the numerical results in Fig. 8 are in agreement with the domain monotonicity property (13), i.e., the curves of $1/(a\mu_{0})$ are ordered according to the fact that a rhombus is inscribed into an ellipse, which in turn is inscribed into a rectangle (of the same aspect ratio $a/b$). In particular, the numerical coefficients in front of $\ln(b/a)$ are $0.52$, $0.59$ and $0.64$ for these three shapes.
For all three shapes, we also observed the asymptotic behavior $\mu_{k}\propto 1/(a\ln(b/a))$ for some other eigenvalues (not shown). It is worth noting that a similar behavior of the principal eigenvalue was obtained for the exterior Steklov problem for elongated prolate spheroids with semiaxes $a$, $a$, $b$ [80]. The explicit construction of a matrix representation of the associated Dirichlet-to-Neumann operator allowed one to show that $\mu_{0,n}\propto 1/(a\ln(b/a))$ as $a\to 0$ for the eigenvalues $\mu_{0,n}$ corresponding to axially symmetric eigenfunctions.
Figure 8: Asymptotic behavior of the principal eigenvalue $\mu_{0}$ for elliptic, rectangular and rhombic patches with $b=1$ and variable $a$. Symbols show $1/(a\mu_{0})$ as a function of $\ln(b/a)$, whereas lines indicate linear fits given in the text.
III.3 Spectral weights
Next, we look at the coefficient $F_{0}$ that represents the relative contribution of the principal eigenmode to the reactive capacitance. For a circular patch, one has $F_{0}\approx 0.97$ [68], i.e., the principal eigenmode provides the dominant contribution. A similar observation was reported for the exterior Steklov problem for tori and two balls [81]. We inspect how the patch anisotropy may affect the coefficient $F_{0}$. Figure 9 shows $F_{0}$ as a function of the aspect ratio $a/b$ for three considered shapes. The first observation is that $F_{0}$ remains the dominant contribution for all three shapes on the considered range $0.01\leq a/b\leq 1$. Curiously, the behavior is quite different between rectangles and elliptic/rhombic patches: the coefficient $F_{0}$ is almost constant for rectangles, whereas it exhibits a very slow decay for ellipses and rhombuses as $a$ decreases. A clarification of this distinction remains an interesting open problem.
Figure 9: Relative contribution $F_{0}$ of the principal eigenmode as a function of $a/b$, for rectangles, ellipses and rhombuses, with $b=1$ and variable $a$.
The empirical observation that $F_{0}$ remains above $0.78$ for the considered domains and aspect ratios raises several geometric questions. Even though $F_{0}$ decreases on the range $0.01\leq a/b\leq 1$, it is yet unclear whether this decay holds in the limit $a\to 0$. Moreover, even if $F_{0}\to 0$ as $a\to 0$, such a slow decay requires to consider extremely thin patches to achieve small $F_{0}$. Is it possible to construct other types of shapes to diminish $F_{0}$? While the actual value of $F_{0}$ does not look relevant, the closeness of $F_{0}$ to $1$ may have important consequences. For instance, the upper and lower bounds (84, 85) are close when $F_{0}\approx 1$, yielding thus an accurate estimate for the principal eigenvalue $\mu_{0}$. Moreover, we showed in Sec. II.4 that the accuracy of the sigmoidal approximation (27) is higher when $F_{0}$ is close to $1$. For these reasons, further analysis and eventual decrease of $F_{0}$ may provide a better understanding of the reactive capacitance.
For such an attempt, we consider a rectangular dumbbell patch formed by two squares that are connected by a long thin rectangular channel. When the channel width goes to 0, the connection progressively disappears, and the dumbbell patch reduces to two disconnected squares. The asymptotic behavior of the eigenvalues and eigenfunctions of the Laplace operator in dumbbell domains was thoroughly investigated in the past (see [103] and references therein). In particular, Laplacian eigenfunctions can be localized in one part of the dumbbell when they cannot “squeeze” through a thin channel, and this behavior strongly depends on the type of the boundary condition (Dirichlet vs Neumann). In contrast, the connectivity of a patch is expected to be much less relevant for the exterior Steklov problem. In fact, even in the disconnected setting, there are long-range interactions between distinct parts of the patch via the kernel $\mathcal{G}(\bm{y},\bm{y}^{\prime})=1/(2\pi|\bm{y}-\bm{y}^{\prime}|)$.
Figure 10 illustrates the first 9 eigenfunctions $\Psi_{k}|_{\Gamma}$ on such a dumbbell patch. The first and most important observation is that the weight $F_{0}\approx 0.56$ of the principal eigenfunction is diminished as compared to that of elliptic, rectangular and rhombic patches. In turn, the contribution of the eigenfunctions $\Psi_{1}$ and $\Psi_{4}$ became more significant, with $F_{1}\approx 0.22$ and $F_{4}\approx 0.20$. Interestingly, we also observe the effect of localization of some eigenfunctions in either of two squares, e.g., $\Psi_{2}$, $\Psi_{3}$ and $\Psi_{6}$ are localized in the larger square, whereas $\Psi_{8}$ is localized in the smaller square. Note that a minor shift of the smaller square breaks the mirror symmetry of the considered dumbbell patch with respect to the horizontal axis, but does not affect the localization of eigenfunctions (not shown).
Figure 10: First 9 Steklov eigenfunctions $\Psi_{k}$ on a rectangular dumbbell patch formed by two squares of lengthsides 11 and $0.7$ and connected by a rectangular channel of length $10$ and width $0.1$ (the associated eigenvalues $\mu_{k}$ and weights $F_{k}$ are shown in the titles). The eigenfunctions were obtained by using a mesh with 1499 triangles and 965 nodes.
III.4 Sigmoidal approximation
We inspect the quality of the explicit approximation $C^{\rm app}(\mu)$ from Eq. (27). For this purpose, we evaluate the maximal relative error of $C^{\rm app}(\mu)$ as compared to $C(\mu)$, over a broad range of $\mu$ from $10^{-2}$ to $10^{2}$. Since the approximation (27) reproduces correctly the leading-order asymptotic behavior of $C(\mu)$ in both limits $\mu\to 0$ and $\mu\to\infty$, this estimate is expected to actually cover the whole range of $\mu$ from 0 to $\infty$. For the considered family of elliptic patches with $a$ from $0.01$ to $1$, the maximal relative error takes the largest value of $4.1\%$ for the circular patch ($a=1$), and then decreases for smaller $a$. A similar trend was observed for rectangular patches, with the maximal relative error of $4.4\%$ corresponding to the quadratic patch.
According to Eq. (34), the relative error of the sigmoidal approximation cannot exceed ${\mathcal{E}}_{\rm max}$. In the case of elliptic patches, one can use the exact relations (90, 143) for $C(\infty)$ and $\mathcal{A}_{\Gamma}$ to get ${\mathcal{E}}_{\rm max}=32/(3\pi^{2})-1\approx 8.1\%$. This conservative bound is twice larger than the actual maximal relative error. For the unit square patch, Eq. (140) yields $\mathcal{A}_{\Gamma}\approx 0.4733$, whereas the electrostatic capacitance was found numerically to be $C(\infty)\approx 0.3667874$ [104] so that ${\mathcal{E}}_{\rm max}\approx 9.1\%$. This bound is again twice larger than the actual maximal relative error.
IV Discussion and conclusion
In this paper, we studied the reactive capacitance $C(\mu)$ for planar patches of arbitrary shape. The spectral representation (14) determines its functional dependence on the reactivity parameter $\mu$, in which the geometric properties of the patch are incorporated through the eigenvalues $\mu_{k}$ and the weights $F_{k}$. Moreover, we developed an efficient numerical tool for computing these geometric parameters for a given patch. This tool was applied to inspect the spectral properties and the resulting reactive capacitances of various patches. We also derived variational formulations, probabilistic interpretations and various bounds for the reactive capacitance. Apart from these theoretical achievements, we obtained several practical results with potential applications in chemical physics that we summarize below.
The first practical result is that the principal eigenmode provides the dominant contribution to the reactive capacitance. On one hand, an important role of the principal eigenmode could be expected; indeed, as $\Psi_{0}$ is the unique eigenfunction that does not change sign on the patch, its projection onto a constant cannot vanish, i.e., $F_{0}>0$. In turn, the other eigenfunctions $\Psi_{k}$ are orthogonal to $\Psi_{0}$ and thus must change sign, so that their projections onto a constant are expected to be smaller. On the other hand, our numerical results indicate that the contribution $F_{0}$ exceeds $96\%$ for all considered rectangular patches, irrespective of their aspect ratio (see Fig. 9). For elliptic and rhombic patches, the situation is different: the relative weight $F_{0}$ is still dominant but it decreases very slowly as the aspect ratio $a/b$ goes to 0. However, even for the value $a/b$ as small as $0.01$, the principal eigenmode provides $88\%$ for the elliptic patch and $78\%$ for the rhombic patch. The dominant role of $F_{0}$ suggests that the associated principal eigenvalue determines the most relevant lengthscale of the problem as $1/\mu_{0}$. Rewriting Eq. (21) as
$C(\infty)=\frac{|\Gamma|}{2\pi}\biggl(\mu_{0}\underbrace{\sum\limits_{k=0}^{\infty}F_{k}}_{=1}+\sum\limits_{k=1}^{\infty}F_{k}(\mu_{k}-\mu_{0})\biggr)$ (97)
and neglecting the second sum, one gets a rough approximation $\mu_{0}\approx 2\pi C(\infty)/|\Gamma|$ (in practice, the right-hand side is the Payne’s upper bound (83)). For instance, for the circular patch of unit radius, this approximation gives $\mu_{0}\approx 4/\pi\approx 1.27$, whereas the actual numerical value is $1.159$, i.e., one gets only $10\%$ relative error. More accurate bounds that involve the constant $\mathcal{A}_{\Gamma}$ were discussed in Sec. II.8. A systematic analysis of the accuracy of these approximations presents an interesting perspective.
Another related open question is to characterize the patches, for which $F_{0}$ is smaller than a given threshold. According to our numerical results, one may probably achieve this condition by taking extremely thin elliptic or rhombic patches, but are there other ways to diminish $F_{0}$? We provided an example of a dumbbell domain, for which the relatively low value $F_{0}\approx 0.56$ could result from the localization phenomenon. However, a rigorous analysis of this phenomenon via local estimates on the eigenfunctions seems to be much more challenging than in the conventional setting of Laplacian eigenfunctions (see [103]). Moreover, one can inspect the contribution of other eigenfunctions and ask whether there exist patches, for which the relative contribution of the principal eigenfunction is not maximal? Even though these questions sound rather abstract, a better understanding of the relation between the shape of the patch and the geometry of its Steklov eigenfunctions may help to design more efficient catalysts with desired reactive properties.
The second practically important result is that the simple sigmoidal approximation $C^{\rm app}(\mu)$ from Eq. (26) turns out to be accurate for a variety of patch shapes that we considered. This numerical observation suggests that, for many practical purposes, it is sufficient to know the electrostatic capacitance $C(\infty)$ and the surface area $|\Gamma|$ of the patch. The additive form of Eq. (26) allows one to interpret $C^{\rm app}(\mu)$ as a consecutive connection of a “diffusion resistance” $1/C(\infty)$ and a “reaction resistance” $2\pi/(\mu|\Gamma|)$, in the language of electrostatics. In fact, as a resistance is the ratio between the applied voltage and the induced electric current, $1/C(\mu)$ plays the role of a resistance in diffusion-reaction problems. The above interpretation was proposed as an empirical way to describe the Laplacian transport, which is decomposed into the first diffusive step and the second reactive step [96, 97, 98, 99]. Its mathematical validation for electrochemical transport was provided in [82], while our work justifies it for steady-state diffusion-controlled reactions on partially reactive flat patches. Moreover, we obtained an upper bound (34) for the relative error of the sigmoidal approximation, which is expressed in terms of $|\Gamma|$, $C(\infty)$, and the constant $\mathcal{A}_{\Gamma}$ defined in Eq. (20). We derived a useful representation (136) for an efficient numerical computation of $\mathcal{A}_{\Gamma}$ for general polygonal patches and provided the exact explicit formulas for elliptic and rectangular patches. The versatile roles of this constant in various contexts were discussed in [81].
Following [90, 81], we introduce the reactive size $L_{\Gamma}=4|\Gamma|/(\pi^{2}C(\infty))$ of a flat patch of arbitrary shape. The numerical prefactor $4/\pi^{2}$ was included for convenience, to ensure that the reactive size of a circular patch is its diameter $2a$. For instance, the reactive size of a quadratic patch with edge length $L$ is $L_{\Gamma}\approx 1.11L$, where we used the numerical value $C(\infty)\approx 0.3667874$ for the capacitance of the unit square [104]. For an elliptic patch with semiaxes $a<b$, one has
$L_{\Gamma}=\frac{4a\,K(\sqrt{1-a^{2}/b^{2}})}{\pi}\,.$ (98)
For elongated elliptic patches with $a\ll b$, one gets $L_{\Gamma}\approx\tfrac{4}{\pi}a\ln(4b/a)$, i.e., the reactive size vanishes as $a\to 0$, as expected.
While this paper was focused on flat patches, most results can be immediately generalized to bounded targets of arbitrary shape with sufficiently smooth boundary in ${\mathbb{R}}^{3}$. In turn, the large-$\mu$ behavior of $C(\mu)$ that originated from the edge singularity, as well as the domain monotonicity of the Steklov eigenvalues, are not valid in general. The efficient numerical tool that we developed here is also not applicable for general targets. In some special cases, the symmetries of the target allow one to employ orthogonal curvilinear coordinates to represent Steklov eigenfunctions on the basis of special functions. This technique was used for prolate and oblate spheroids in [80], as well as for a pair of balls and a torus in [81]. For instance, the spectral expansion (14), the role of the principal weight $F_{0}$, and the accuracy of the sigmoidal approximation (27) were discussed for these targets in [81]. However, the numerical computation of Steklov eigenmodes for an arbitrarily-shaped target requires the knowledge of the Green’s function $G(\bm{y},\bm{y}^{\prime})$ in the exterior of the target, as well as a suitable boundary integral method. A further development in this direction presents an interesting perspective. More generally, the proposed theoretical description of the reactive capacitance opens promising ways for modeling diffusion-reaction phenomena in complex media and for solving shape-optimization problems in chemical engineering and other disciplines.
Acknowledgements.
The authors thank Prof. N. Nigam and Prof. I. Polterovich for fruitful discussions. D. S. G. acknowledges the Simons Foundation for supporting his sabbatical sojourn in 2024 at the CRM, University of Montréal, Canada, as well as the Alexander von Humboldt Foundation for support within a Bessel Prize award.
Appendix A Numerical approach
This Appendix describes the practical implementation of our numerical approach for solving the spectral problems (3) and (23).
A.1 Finite-element method
We employ a finite-element method to solve the integral eigenvalue problems numerically. The patch $\Gamma$ is covered by a triangular mesh, composed of $N_{t}$ triangles denoted as $T_{j}$, $j=1,\ldots,N_{t}$. We denote three vertices of the $j$-th triangle as $P_{jk}=(P_{jk}^{x},P_{jk}^{y})$, with $k=1,2,3$, and the total number of vertices is $N_{p}$. For convenience of notations, we allow the index $k$ to take values $4$ and $5$ by setting $P_{j4}=P_{j1}$ and $P_{j5}=P_{j2}$. We also introduce the unit tangent vector $\bm{\tau}_{jk}$ and the unit normal vector $\bm{n}_{jk}$ to the edge $(P_{jk},P_{j(k+1)})$ (see Fig. 11):
$\bm{\tau}_{jk}$ $=\frac{P_{j(k+1)}-P_{jk}}{|P_{j(k+1)}-P_{jk}|}\,$, (99)
$\bm{n}_{jk}$ $=\pm(\bm{\tau}_{jk}^{y},-\bm{\tau}_{jk}^{x})^{\dagger}$ (100)
(the sign $\pm$ is adjusted to ensure that $\bm{n}_{jk}$ is oriented outward the triangle).
Figure 11: Triangle $T_{j}$ of a mesh, its vertices $P_{jk}$ and normal and tangent vectors $\bm{n}_{jk}$ and $\bm{\tau}_{jk}$.
We employ linear basis functions to represent eigenfunctions on the mesh in order to achieve higher accuracy as compared to piecewise-constant basis functions. For the sake of clarify, we start with barycentric basis functions $\{\varphi_{jk}\}$, i.e., $\varphi_{jk}(\bm{y})$ is a linear function on the $j$-th triangle $T_{j}$ (and 0 otherwise) such that $\varphi_{jk}(P_{jk^{\prime}})=\delta_{k,k^{\prime}}$ for $k,k^{\prime}\in\{1,2,3\}$. They form so-called discontinuous Galerkin or “broken P1” space, to be distinguished from nodal basis (or hat) functions that we will reconstruct at the end. The barycentric basis functions can be written explicitly as
$\varphi_{jk}(\bm{y})=1-\langle\bm{y}-P_{jk},\bm{v}_{jk}\rangle,$ (101)
where $\langle\cdot,\cdot\rangle$ denotes the scalar product, and
[…]
ensures that $\varphi_{jk}(P_{j(k+1)})=\varphi_{jk}(P_{j(k+2)})=0$ (with $|T_{j}|$ being the surface area of the $j$-th triangle). One can easily check that
$\bm{v}_{jk}=\frac{|P_{j(k+2)}-P_{j(k+1)}|}{2|T_{j}|}\bm{n}_{j(k+1)}\,.$ (103)
We use the double index $jk$ to enumerate the basis functions, as well as matrix elements constructed with the help of these functions. An eigenfunction $\Psi(\bm{y})$ is searched as a linear superposition:
$\Psi(\bm{y})=\sum\limits_{j=1}^{N_{t}}\sum\limits_{k=1}^{3}V_{jk}\varphi_{jk}(\bm{y}),$ (104)
with unknown coefficients $V_{jk}$. Substituting this superposition into the integral equation, multiplying it by another basis function $\varphi_{j^{\prime}k^{\prime}}$ and integrating over $\Gamma$, we get the system of $3N_{t}$ linear equations
$\sum\limits_{j=1}^{N_{t}}\sum\limits_{k=1}^{3}V_{jk}G_{jk,j^{\prime}k^{\prime}}=\frac{1}{\mu}\sum\limits_{j=1}^{N_{t}}\sum\limits_{k=1}^{3}V_{jk}M_{jk,j^{\prime}k^{\prime}},$ (105)
where
$M_{jk,j^{\prime}k^{\prime}}$ $=\int\limits_{\Gamma}\varphi_{jk}\varphi_{j^{\prime}k^{\prime}}d\bm{y}=\delta_{j,j^{\prime}}\int\limits_{T_{j}}\varphi_{jk}\varphi_{j^{\prime}k^{\prime}}\,d\bm{y}$
$=\delta_{j,j^{\prime}}\frac{|T_{j}|}{12}(1+\delta_{k,k^{\prime}})$ (106)
is the standard mass matrix, and
$G_{jk,j^{\prime}k^{\prime}}=\int\limits_{\Gamma}d\bm{x}\,\varphi_{j^{\prime}k^{\prime}}(\bm{x})\int\limits_{\Gamma}\varphi_{jk}(\bm{y})\,\mathcal{G}(\bm{x},\bm{y})d\bm{y}$ (107)
is the matrix representation of the kernel $\mathcal{G}(\bm{x},\bm{y})$ given by Eq. (10).
Up to now, this was a standard procedure for an arbitrary kernel $\mathcal{G}(\bm{x},\bm{y})$. In the following, we rely on the explicit form of the kernel and the flat geometry of the patch to obtain a semi-analytical representation of the matrix elements $G_{jk,j^{\prime}k^{\prime}}$. In fact, we aim to evaluate exactly the integral
$Q_{jk}(\bm{x})=\int\limits_{\Gamma}\varphi_{jk}(\bm{y})\mathcal{G}(\bm{x},\bm{y})\,d\bm{y}.$ (108)
Using Eq. (101), we get
$Q_{jk}(\bm{x})=\frac{F_{j}(\bm{x})}{2\pi}\bigl(1+\langle P_{jk},\bm{v}_{jk}\rangle\bigr)-\frac{H_{j}(\bm{x},\bm{v}_{jk})}{2\pi}\,$, (109)
where
$F_{j}(\bm{x})$ $=\int\limits_{T_{j}}\frac{d\bm{y}}{|\bm{x}-\bm{y}|}\,$, (110)
$H_{j}(\bm{x},\bm{v})$ $=\int\limits_{T_{j}}\frac{\langle\bm{v},\bm{y}\rangle\,d\bm{y}}{|\bm{x}-\bm{y}|}\,.$ (111)
For the first integral, we use the identity $\Delta_{\bm{y}}|\bm{x}-\bm{y}|=1/|\bm{x}-\bm{y}|$ to get by the divergence theorem
$F_{j}(\bm{x})$ $=\int\limits_{\partial T_{j}}\langle\bm{n}_{\bm{y}},\nabla_{\bm{y}}|\bm{x}-\bm{y}|\rangle dl_{\bm{y}}$
$=\sum\limits_{k=1}^{3}\int\limits_{\partial T_{jk}}\frac{\langle\bm{n}_{\bm{y}},(\bm{y}-\bm{x})\rangle}{|\bm{y}-\bm{x}|}dl_{\bm{y}},$ (112)
where $\partial T_{jk}$ denotes the edge of the $j$-th triangle between vertices $P_{jk}$ and $P_{j(k+1)}$ (here $\Delta_{\bm{y}}$ is the two-dimensional Laplace operator). Using the parameterization $\bm{y}(t)=P_{j,k}+(P_{j(k+1)}-P_{jk})t$ on this edge (with $t\in(0,1)$ and $dl_{\bm{y}}=|\bm{y}^{\prime}(t)|dt$), we have
$F_{j}(\bm{x})=\sum\limits_{k=1}^{3}|P_{j(k+1)}-P_{jk}|\int\limits_{0}^{1}\frac{\langle\bm{n}_{jk},(\bm{y}(t)-\bm{x})\rangle}{|\bm{y}(t)-\bm{x}|}dt,$ (113)
where $\bm{n}_{jk}$ is the outward unit normal vector to the edge $(P_{jk},P_{j(k+1)})$. After lengthy but elementary computation of the last integral, we get
$F_{j}(\bm{x})=\sum\limits_{k=1}^{3}\langle\bm{n}_{jk},R_{jk}\rangle\,I_{jk},$ (114)
where $R_{jk}=P_{jk}-\bm{x}$, and
$I_{jk}=\ln\left(\frac{\langle R_{j(k+1)},\bm{\tau}_{jk}\rangle+|R_{j(k+1)}|}{\langle R_{jk},\bm{\tau}_{jk}\rangle+|R_{jk}|}\right),$ (115)
with the unit tangent vector $\bm{\tau}_{jk}$ given by Eq. (99).
For the integral (111), we use $\langle\bm{v},\nabla_{\bm{y}}|\bm{x}-\bm{y}|\rangle=\langle\bm{v},(\bm{y}-\bm{x})\rangle/|\bm{y}-\bm{x}|$ and the divergence theorem to get
$H_{j}(\bm{x},\bm{v})=\langle\bm{v},\bm{x}\rangle F_{j}(\bm{x})+\int\limits_{\partial T_{j}}|\bm{x}-\bm{y}|\langle\bm{v},\bm{n}_{\bm{y}}\rangle\,d\bm{y}\,.$ (116)
As previously, we use the parameterization over each edge to get
$H_{j}(\bm{x},\bm{v})=\langle\bm{v},\bm{x}\rangle F_{j}(\bm{x})+\sum\limits_{k=1}^{3}\langle\bm{v},\bm{n}_{jk}\rangle J_{jk},$ (117)
where
$J_{jk}$ $=\int\limits_{\partial T_{jk}}|\bm{x}-\bm{y}|\,dl_{\bm{y}}=\frac{I_{jk}}{2}\bigl(|R_{jk}|^{2}-\langle R_{jk},\bm{\tau}_{jk}\rangle^{2}\bigr)$
$+\frac{1}{2}\biggl(\langle R_{j(k+1)},\bm{\tau}_{jk}\rangle|R_{j(k+1)}|-\langle R_{jk},\bm{\tau}_{jk}\rangle|R_{jk}|\biggr).$ (118)
As the function $Q_{jk}(\bm{x})$ is known exactly, the matrix elements can be found as
$G_{jk,j^{\prime}k^{\prime}}=\int\limits_{T_{j^{\prime}}}\varphi_{j^{\prime}k^{\prime}}(\bm{x})\,Q_{jk}(\bm{x})\,d\bm{x}.$ (119)
The simplest quadrature employs the barycentric rule. However, one of the major numerical difficulties for solving the spectral problem (11) is that the matrix $\bf G$} is not sparse that significantly limits its size and thus requires using relatively coarse meshing. In order to improve the accuracy, it is therefore convenient to apply more accurate quadratures when integrating over triangles. We employ Dunavant 7-point quadrature rule [105]: for a smooth enough function $f(\bm{y})$, we set
$\int\limits_{T_{j}}f(\bm{y})\,d\bm{y}\approx|T_{j}|\sum\limits_{i=1}^{7}w_{i}\,f\bigl(\nu_{i1}P_{j1}+\nu_{i2}P_{j2}+\nu_{i3}P_{j3}\bigr),$ (120)
where
$w_{1}$ $=0.225,\quad\nu_{11}=\nu_{12}=\nu_{13}=1/3,$
$w_{2}$ $=0.132394,~~\nu_{21}=0.059716,~~\nu_{22}=\nu_{23}=0.470142,$
$w_{5}$ $=0.125939,~~\nu_{51}=0.797427,~~\nu_{52}=\nu_{53}=0.101287,$
whereas $w_{3}=w_{4}=w_{2}$ with $\nu_{3i}$ and $\nu_{4i}$ being obtained by permutations from $\nu_{2i}$, and $w_{6}=w_{7}=w_{5}$ with $\nu_{6i}$ and $\nu_{7i}$ being obtained by permutations from $\nu_{5i}$. This quadrature is known to be exact for polynomials up to degree 5.
Once the matrix elements $G_{jk,j^{\prime}k^{\prime}}$ are found, one can solve numerically the generalized eigenvalue problem for square matrices $\bf G$} and $\bf M$}, that allows one to approximate some Steklov eigenvalues and eigenfunctions. This final step can be further improved by transforming the above representation in terms of barycentric basis functions $\{\varphi_{jk}\}$, associated with triangles, into more convenient nodal basis (or hat) functions $\{\phi_{i}\}$, associated to the nodes of the mesh. For each node $P_{i}$ (with $i=1,2,\ldots,N_{p}$), $\phi_{i}(\bm{x})$ is a linear function on the triangles that contain $P_{i}$ (and 0 otherwise) such that $\phi_{i}(P_{i})=1$ and 0 at all other nodes. In fact, on the $j$-th triangle containing the vertex $P_{i}$, $\phi_{i}$ is simply the barycentric function $\varphi_{jk}$, with the index $k$ such that $P_{jk}=P_{i}$. As a consequence, one can regroup the earlier defined matrix elements $G_{jk,j^{\prime}k^{\prime}}$ and $M_{jk,j^{\prime}k^{\prime}}$ into new matrix elements $\bar{G}_{i,i^{\prime}}$ and $\bar{M}_{i,i^{\prime}}$ as
$\bar{G}_{i,i^{\prime}}$ $=\sum\limits_{j,k:\atop P_{jk}=P_{i}}\sum\limits_{j^{\prime},k^{\prime}:\atop P_{j^{\prime}k^{\prime}}=P_{i^{\prime}}}G_{jk,j^{\prime}k^{\prime}}\quad(i,i^{\prime}=1,\ldots,N_{p}),$ (121)
$\bar{M}_{i,i^{\prime}}$ $=\sum\limits_{j,k:\atop P_{jk}=P_{i}}\sum\limits_{j^{\prime},k^{\prime}:\atop P_{j^{\prime}k^{\prime}}=P_{i^{\prime}}}M_{jk,j^{\prime}k^{\prime}}\quad(i,i^{\prime}=1,\ldots,N_{p}).$ (122)
In this way, one has to solve numerically the generalized eigenvalue problem for square matrices $\bar{{\bf G}}$ and $\bar{{\bf M}}$:
$\bar{{\bf G}}{\bf V}=\lambda\bar{{\bf M}}{\bf V}.$ (123)
Let us denote the eigenvalues and eigenvectors of this problem as $\lambda_{k}$ and ${\bf V}_{k}$ and enumerate the eigenvalues to form a decreasing sequence: $\lambda_{0}\geq\lambda_{1}\geq\lambda_{2}\geq\cdots$. One sees that $1/\lambda_{k}$ is an approximation of the $k$-th eigenvalue $\mu_{k}$, whereas the components of the vector ${\bf V}_{k}$ are the coefficients of an approximate expansion of the $k$-th eigenfunction $\Psi_{k}$ on the nodal basis functions:
$\Psi_{k}(\bm{y})=\sum\limits_{i=1}^{N_{p}}[{\bf V}_{k}]_{i}\,\phi_{i}(\bm{y}).$ (124)
Moreover, the value of $\Psi_{k}$ on a node $P_{i}$ is simply ${\bf V}_{k}]_{i}$.
Note that the normalization of eigenfunctions is fixed by requiring
$1=\int\limits_{\Gamma}\Psi_{k}^{2}(\bm{y})\,d\bm{y}=\sum\limits_{i,i^{\prime}=1}^{N_{p}}[{\bf V}_{k}]_{i}[{\bf V}_{k}]_{i^{\prime}}\underbrace{\int\limits_{\Gamma}\phi_{i}\,\phi_{i^{\prime}}\,d\bm{y}}_{=\bar{M}_{i,i^{\prime}}}.$ (125)
We also compute the projection of $\Psi_{k}$ onto a constant:
$d_{k}=\int\limits_{\Gamma}\Psi_{k}(\bm{y})\,d\bm{y}=\sum\limits_{i=1}^{N_{p}}[{\bf V}_{k}]_{i}\underbrace{\int\limits_{\Gamma}\phi_{i}(\bm{y})\,d\bm{y}}_{=\sum\limits_{j,k^{\prime}:P_{jk^{\prime}}=P_{i}}|T_{j}|/3},$ (126)
from which the weights follow as $F_{k}=d_{k}^{2}/|\Gamma|$. We outline that the representation in terms of nodal basis functions has two practical advantages: (i) the matrices $\bar{{\bf G}}$ and $\bar{{\bf M}}$ of size $N_{p}\times N_{p}$ are much smaller than $\bf G$} and $\bf M$} (of size $3N_{t}\times 3N_{t}$) that allows for their faster numerical diagonalization; and (ii) the continuous nature of nodal basis functions yields a smoother representation of eigenfunctions. We employ this method throughout the manuscript.
A.2 Solving the alternative Steklov problem
An equivalent reformulation of the spectral problem (23) was discussed in [83, 66]. To ensure that all eigenfunctions $\Psi_{k}^{N}$ with $k=1,2,\ldots$ are orthogonal to the constant function $\Psi_{0}^{N}=1/\sqrt{|\Gamma|}$, the kernel $\mathcal{G}(\bm{y},\bm{y}^{\prime})$ can be replaced by
$\mathcal{G}^{N}(\bm{y},\bm{y}^{\prime})=\mathcal{G}(\bm{y},\bm{y}^{\prime})-\frac{\omega_{\Gamma}(\bm{y})+\omega_{\Gamma}(\bm{y}^{\prime})}{|\Gamma|}+\mathcal{A}_{\Gamma},$ (127)
where $\omega_{\Gamma}(\bm{y})$ and $\mathcal{A}_{\Gamma}$ were defined in Eqs. (19, 20). In this way, the integral of $\mathcal{G}^{N}(\bm{y},\bm{y}^{\prime})$ over $\bm{y}\in\Gamma$ vanishes so that a constant function is an eigenfunction of the associated integral operator. It was shown in [66] that the eigenfunctions $\Psi_{k}^{N}$ satisfy for $k=1,2,\ldots$:
$\int\limits_{\Gamma}\mathcal{G}^{N}(\bm{y},\bm{y}^{\prime})\Psi_{k}^{N}(\bm{y}^{\prime})d\bm{y}^{\prime}=\frac{1}{\mu_{k}^{N}}\Psi_{k}^{N}(\bm{y})\qquad(\bm{y}\in\Gamma).$ (128)
In other words, one can search for the eigenpairs $\{\mu_{k}^{N},\Psi_{k}^{N}\}$ (with $k\geq 1$) of the spectral problem (23) by solving the eigenvalue problem (128). The missing eigenpair $\mu_{0}^{N}=0$ and $\psi_{0}^{N}=1/\sqrt{|\Gamma|}$ can be added manually.
The numerical technique from Sec. A.1 is applicable for solving the Steklov problem (23), which is equivalently described by the kernel $\mathcal{G}^{N}(\bm{y},\bm{y}^{\prime})$ from Eq. (127). In fact, we need to find
$Q_{jk}^{N}(\bm{x})$ $=\int\limits_{\Gamma}\varphi_{jk}(\bm{y})\mathcal{G}^{N}(\bm{x},\bm{y})d\bm{y}$
$\approx Q_{jk}(\bm{x})-\biggl[\frac{\omega_{\Gamma}(\bm{x})+\omega_{\Gamma}(\bm{x}_{j})}{|\Gamma|}-\mathcal{A}_{\Gamma}\biggr]\frac{|T_{j}|}{3}\,$, (129)
where $\bm{x}_{j}$ is the barycenter of $T_{j}$, and we used the barycentric approximation $\int\nolimits_{T_{j}}\varphi_{jk}(\bm{y})\,\omega_{\Gamma}(\bm{y})d\bm{y}\approx\omega_{\Gamma}(\bm{x}_{j})\int\nolimits_{T_{j}}\varphi_{jk}(\bm{y})\,d\bm{y}$ (it can further be improved via Dunavant 7-point quadrature). Using the approximations
$\omega_{\Gamma}(\bm{x})=\frac{1}{2\pi}\sum\limits_{j=1}^{N}F_{j}(\bm{x})\,,\quad\mathcal{A}_{\Gamma}\approx\frac{1}{|\Gamma|}\sum\limits_{j=1}^{N}|T_{j}|w(\bm{x}_{j}),$ (130)
we get
$G_{jk,j^{\prime}k^{\prime}}^{N}$ $\approx\frac{|T_{j^{\prime}}|}{3}Q_{jk}^{N}(\bm{x}_{j^{\prime}})$ (131)
$\approx G_{jk,j^{\prime}k^{\prime}}-\frac{|T_{j}|\,|T_{j^{\prime}}|}{9}\biggl[\frac{\omega_{\Gamma}(\bm{x}_{j})+\omega_{\Gamma}(\bm{x}_{j^{\prime}})}{|\Gamma|}-\mathcal{A}_{\Gamma}\biggr]\,$,
from which the matrix elements $\bar{G}_{i,i^{\prime}}^{N}$ are deduced via Eq. (121). As previously, one searches for eigenpairs $\{\lambda_{k}^{N},{\bf V}_{k}^{N}\}$ solving the generalized eigenvalue problem $\bar{{\bf G}}^{N}{\bf V}_{k}^{N}=\lambda_{k}^{N}\bar{{\bf M}}{\bf V}_{k}^{N}$, enumerated by $k=1,2,\ldots$ (here the index starts from $1$, given that $\mu_{0}^{N}=0$).
Once the eigenfunction $\Psi_{k}^{N}(\bm{y})$ is constructed on the patch, it can be easily extended into the upper half-space ${\mathbb{R}}^{3}_{+}$. For this purpose, one can multiply Eq. (23) by the Green’s function $G(\bm{y},\bm{y}^{\prime})$, multiply Eq. (6a) by $\Psi_{k}^{N}(\bm{y}^{\prime})$, subtract them, integrate over $\bm{y}^{\prime}\in{\mathbb{R}}^{3}_{+}$, use the Green’s formula with boundary conditions and behavior at infinity to get
$\Psi_{k}^{N}(\bm{y})=\Psi_{k}^{N}(\infty)+\mu_{k}^{N}\int\limits_{\Gamma}G(\bm{y},\bm{y}^{\prime})\Psi_{k}^{N}(\bm{y}^{\prime})d\bm{y},$ (132)
where $G(\bm{y},\bm{y}^{\prime})$ is given by the explicit formula (7). In turn, the value at infinity, $\Psi_{k}^{N}(\infty)$, can be determined by expressing $G(\bm{y},\bm{y}^{\prime})$ from Eq. (127) and using Eq. (128):
$\Psi_{k}^{N}(\infty)=-\frac{\mu_{k}^{N}}{|\Gamma|}\int\limits_{\Gamma}\omega_{\Gamma}(\bm{y})\,\Psi_{k}^{N}(\bm{y})d\bm{y}.$ (133)
A.3 Validation
To check the accuracy of our numerical method, we compute the Steklov eigenvalues for the circular patch of unit radius. This geometric setting can be considered as the limit $c\to 0$ of an oblate spheroid with semi-axes $1,1,c$, for which an efficient numerical method for solving Steklov problems was developed in [80]. This method employs the oblate spheroidal coordinates to achieve an accurate matrix representation of the associated Dirichlet-to-Neumann operator. Its numerical diagonalization yields very accurate eigenvalues, which will be referred to as “exact”.
Table 1 reports the first ten eigenvalues for both versions of the Steklov problem (see also Fig. 2 for eigenfunctions $\Psi_{k}$). Two different meshes were used to illustrate the role of the mesh size on the accuracy. The relative error of the presented eigenvalues does not exceed $0.1\%$ even for the coarser mesh.
The rotational symmetry of the patch implies that an eigenfunction on $\Gamma$ can be represented in polar coordinates $(r,\phi)$ as $e^{im\phi}v_{m,n}(r)$, with integer indices $m$ and $n$. Moreover, the structure of radial functions $v_{m,n}(r)$ is inherited from spherical harmonics $Y_{m,n}$ (see [80] for more details); in particular, one can set $n=0,1,2,\cdots$ and $m=-n,-n+1,\cdots,n$, as shown in the bottom row of Table 1. Note that the eigenvalues with $(\pm m)n$ coincide. As discussed in [80], the eigenfunctions satisfying Neumann boundary condition (3c) correspond to even $m+n$. Finally, as the difference $\mathcal{G}(\bm{x},\bm{y})-\mathcal{G}^{N}(\bm{x},\bm{y})$ is a radial function (given that $\omega_{\Gamma}(\bm{y})$ is a function of $|\bm{y}|$), only axially symmetric eigenfunctions with $m=0$ differ between the two versions of the Steklov problem. This is clearly seen from Table 1. Note also that our approach does not recover the eigenvalue $\mu_{0}^{N}=0$ that corresponds to a constant eigenfunction.
$N_{p}$ $N_{t}$ $k$ 0 1 2 3 4 5 6 7 8 9
375 688 D 1.1588 2.7573 2.7573 4.1252 4.1252 4.3209 5.4053 5.4053 5.8984 5.8984
553 1032 D 1.1585 2.7565 2.7565 4.1240 4.1240 4.3196 5.4036 5.4036 5.8963 5.8963
exact D 1.1578 2.7548 2.7548 4.1214 4.1214 4.3169 5.4003 5.4003 5.8924 5.8924
375 688 N 2.7573 2.7573 4.1252 4.1252 4.1239 5.4053 5.4053 5.8984 5.8984
553 1032 N 2.7565 2.7565 4.1240 4.1240 4.1230 5.4036 5.4036 5.8962 5.8963
exact N 0 2.7573 2.7573 4.1252 4.1252 4.1213 5.4003 5.4003 5.8924 5.8924
mnmn $00$ (1)1(-1)1 $11$ (2)2(-2)2 $22$ $02$ (3)3(-3)3 $33$ (1)3(-1)3 $13$
Table 1: Comparison of the Steklov eigenvalues for the circular patch of unit radius between our FEM method and semi-analytical computation via oblate spheroidal coordinates in [80, 66], referred to as “exact”. The number of vertices ($N_{p}$) and the number of triangle ($N_{t}$) are indicated, whereas the letters D and N distinguish respectively the spectral problems (3) and (23). The bottom line indicates an alternative enumeration of the eigenpairs according to the symmetries of the eigenfunctions. For the “Neumann version”, the eigenvalue $4.1213$ and its numerical estimates were placed into the column with $k=5$ to respect the radial symmetry $mn=02$ of the associated eigenfunction, despite the artificially induced error in ordering of the eigenvalues.
Appendix B Computation of the constant $\mathcal{A}_{\Gamma}$
The function $\omega_{\Gamma}(\bm{x})$ from Eq. (19) and the constant $\mathcal{A}_{\Gamma}$ from Eq. (20) play an important role in the reformulation of the spectral problem (23); see also [66] for various applications of these quantities. For a circular patch, their exact forms were obtained in [66]:
$\omega_{\Gamma}(\bm{x})=\frac{2}{\pi}E(|\bm{x}|),\qquad\mathcal{A}_{\Gamma}=\frac{8}{3\pi^{2}}\,$, (134)
where $E(z)$ is the complete elliptic integral of the second kind. In this Appendix, we discuss an alternative computation of these quantities for other domains, beyond the basic FEM representation (130).
B.1 General patch
For any triangular patch, the function $\omega_{\Gamma}(\bm{x})$ is simply proportional to the function $F_{j}(\bm{x})$ defined by Eq. (110), for which the representation (114) provides the exact solution. Moreover, the exact representation (114) can be generalized to any polygon. As a consequence, this representation allows one to compute exactly the function $w_{\Gamma}(\bm{x})$ from Eq. (19) for any polygonal patch, without resorting to triangulations:
$\omega_{\Gamma}(\bm{x})=\frac{1}{2\pi}\sum\limits_{k}\langle\bm{n}_{k},(P_{k}-\bm{x})\rangle I_{k},$ (135)
where $P_{k}$ are the vertices of the polygon, $\bm{n}_{k}$ is the unit normal vector to the $k$-th edge between vertices $P_{k}$ and $P_{k+1}$ ($\bm{n}_{k}$ is oriented outward the polygon), $I_{k}$ is given by Eq. (115), and the sum is over all edges. In turn, the constant $\mathcal{A}_{\Gamma}$ can be written as follows
$\mathcal{A}_{\Gamma}$ $=\frac{1}{2\pi|\Gamma|^{2}}\int\limits_{\partial\Gamma}dl_{\bm{y}}\langle\bm{n}_{\bm{y}},\nabla_{\bm{y}}\rangle\int\limits_{\Gamma}|\bm{x}-\bm{y}|\,d\bm{x}$
$=\frac{1}{2\pi|\Gamma|^{2}}\int\limits_{\partial\Gamma}dl_{\bm{y}}\langle\bm{n}_{\bm{y}},\nabla_{\bm{y}}\rangle\int\limits_{\partial\Gamma}\langle\bm{n}_{\bm{x}},\nabla_{\bm{x}}|\bm{x}-\bm{y}|^{3}/9\rangle dl_{\bm{x}},$
where we used that $\Delta_{\bm{x}}|\bm{x}-\bm{y}|^{3}/9=|\bm{x}-\bm{y}|$. As a consequence, we obtain
$\mathcal{A}_{\Gamma}$ $=\frac{1}{6\pi|\Gamma|^{2}}\int\limits_{\partial\Gamma}dl_{\bm{y}}\langle\bm{n}_{\bm{y}},\nabla_{\bm{y}}\rangle\int\limits_{\partial\Gamma}\langle\bm{n}_{\bm{x}},\nabla_{\bm{x}}|\bm{x}-\bm{y}|(\bm{x}-\bm{y})\rangle dl_{\bm{x}}$
$=-\frac{1}{6\pi|\Gamma|^{2}}\int\limits_{\partial\Gamma\times\partial\Gamma}\biggl(\langle\bm{n}_{\bm{x}},\bm{n}_{\bm{y}}\rangle|\bm{x}-\bm{y}|$
$+\frac{\langle\bm{n}_{\bm{x}},\bm{x}-\bm{y}\rangle\,\langle\bm{n}_{\bm{y}},\bm{x}-\bm{y}\rangle}{|\bm{x}-\bm{y}|}\biggr)dl_{\bm{x}}dl_{\bm{y}}.$ (136)
This formula opens a purely geometric way to access $\mathcal{A}_{\Gamma}$, allowing one to skip the computation of $\omega_{\Gamma}(\bm{x})$ and its integration. We recall that the constant $\mathcal{A}_{\Gamma}$ determines via Eq. (20) the coefficient $c_{2}$ of the Taylor expansion (17) of the reactive capacitance.
B.2 Rectangular patch
For a rectangle $(-a,a)\times(-b,b)$, Eq. (135) yields
$\omega_{\Gamma}(x,y)=\frac{1}{2\pi}$ (137)
$\times\biggl\{(b+y)\ln\biggl(\frac{a-x+\sqrt{(a-x)^{2}+(b+y)^{2}}}{-a-x+\sqrt{(a+x)^{2}+(b+y)^{2}}}\biggr)$
$+(a-x)\ln\biggl(\frac{b-y+\sqrt{(a-x)^{2}+(b-y)^{2}}}{-b-y+\sqrt{(a-x)^{2}+(b+y)^{2}}}\biggr)$
$+(b-y)\ln\biggl(\frac{a+x+\sqrt{(a+x)^{2}+(b-y)^{2}}}{-a+x+\sqrt{(a-x)^{2}+(b-y)^{2}}}\biggr)$
$+(a+x)\ln\biggl(\frac{b+y+\sqrt{(a+x)^{2}+(b+y)^{2}}}{-b+y+\sqrt{(a+x)^{2}+(b-y)^{2}}}\biggr)\biggr\}.$
After a lengthy but elementary integration of this expression over the rectangle $(-a,a)\times(-b,b)$, we deduce the associated constant $\mathcal{A}_{\Gamma}$:
$\mathcal{A}_{\Gamma}=\frac{a\,\eta(b/a)}{\pi b^{2}}+\frac{b\,\eta(a/b)}{\pi a^{2}}\,$, (138)
where
$\eta(y)$ $=\frac{1+y^{3}-(1+y^{2})^{3/2}}{3}+\frac{\sqrt{1+y^{2}}-1}{2}$
$+\frac{y^{2}}{4}\ln\biggl(\frac{1+\sqrt{1+y^{2}}}{-1+\sqrt{1+y^{2}}}\biggr).$ (139)
For instance, we get for the square of edge $2b$
$\mathcal{A}_{\Gamma}=\frac{2\eta(1)}{\pi b}\,$, (140)
with $\eta(1)=\tfrac{1}{6}(1-\sqrt{2})+\tfrac{1}{2}\ln(1+\sqrt{2})\approx 0.3717$.
In the limit $a\to 0$, we use the asymptotic behavior of the function $\eta(y)$, namely, $\eta(y)\simeq y/2+O(1)$ as $y\to\infty$ and $\eta(y)\simeq y^{2}(-1+2\ln(2/y))/4+O(y^{3})$ as $y\to 0$, to obtain
$\mathcal{A}_{\Gamma}\simeq\frac{1}{\pi b}\biggl(\frac{1+2\ln 2}{4}+\frac{1}{2}\ln(b/a)\biggr)+O(a/b).$ (141)
In fact, as the rectangular patch shrinks to an interval, which is inaccessible to reflected Brownian motion, the constant $\mathcal{A}_{\Gamma}$ diverges logarithmically.
B.3 Elliptic patch
For an elliptic patch with semi-axes $a$ and $b$, we employ the parameterization $\bm{x}_{1}(t_{1})=(a\cos t_{1},b\sin t_{1})^{\dagger}$ so that $c(t_{1})=|\tfrac{d}{dt_{1}}\bm{x}_{1}(t_{1})|=\sqrt{a^{2}\sin^{2}t_{1}+b^{2}\cos^{2}t_{1}}$ and $\bm{n}_{\bm{x}_{1}}=(b\cos t_{1},a\sin t_{1})^{\dagger}/c(t_{1})$ and similar for $\bm{x}_{2}(t_{2})$. As a consequence, Eq. (136) reads
$\mathcal{A}_{\Gamma}=\frac{-1}{6\pi^{3}b}\int\limits_{0}^{2\pi}dt_{1}\int\limits_{0}^{2\pi}dt_{2}\biggl\{(\gamma^{-2}\cos t_{1}\cos t_{2}+\sin t_{1}\sin t_{2})$
$\times D(t_{1},t_{2})-\frac{[1-(\cos t_{1}\cos t_{2}+\sin t_{1}\sin t_{2})]^{2}}{D(t_{1},t_{2})}\biggr\},$ (142)
where $\gamma=a/b$ and
$D(t_{1},t_{2})=\sqrt{\gamma^{2}(\cos t_{1}-\cos t_{2})^{2}+(\sin t_{1}-\sin t_{2})^{2}}\,.$
To proceed, one can split the integration domain $(0,2\pi)\times(0,2\pi)$ into four twice smaller squares and use the symmetry of sine and cosine functions to reduce the integration domain to $(0,\pi)\times(0,\pi)$. By introducing the new integration variables $t_{\pm}=(t_{1}\pm t_{2})/2$, these integrals can be reduced to the complete elliptic functions. Skipping these tedious lengthy computations, we just provide the final result, which takes a particularly simple form:
$\mathcal{A}_{\Gamma}=\frac{16}{3\pi^{3}b}K\bigl(\sqrt{1-a^{2}/b^{2}}\bigr),$ (143)
where $K(z)$ is defined by Eq. (75). In the particular case of a circular patch ($a=b$), we use $K(0)=\pi/2$ to retrieve the result (134). In turn, in the limit $a\to 0$, we use the asymptotic relation (91) to get
$\mathcal{A}_{\Gamma}\simeq\frac{16}{3\pi^{3}b}\ln(4b/a)\quad(a\ll b).$ (144)
Substituting Eq. (143) into our upper bound (86), we get for any $0<a\leq b$:
$\mu_{0}\leq\frac{3\pi^{2}}{16aK(\sqrt{1-a^{2}/b^{2}})}\,.$ (145)
It is instructive to compare this inequality with the Payne’s upper bound (83). Substituting Eq. (90) for the electrostatic capacitance, one has
$\mu_{0}\leq\frac{2}{aK(\sqrt{1-a^{2}/b^{2}})}\,.$ (146)
Remarkably, both bounds have the same functional form and differ only by the numerical prefactors: $3\pi^{2}/16\approx 1.85$ in Eq. (145) and $2$ in Eq. (146). One sees that our upper bound is slightly more accurate for any $0<a\leq b$.
B.4 Circular annulus
Finally, we consider a circular annulus of radii $R_{1}<R_{2}$. As the boundary of the annulus is split into two circles, the integral in Eq. (136) has four contributions when $\bm{x}_{1}$ and $\bm{x}_{2}$ run over the inner/outer circles:
$\mathcal{A}_{\Gamma}=\mathcal{A}^{ii}_{\Gamma}+\mathcal{A}^{oo}_{\Gamma}+\mathcal{A}^{io}_{\Gamma}+\mathcal{A}^{oi}_{\Gamma},$
where superscripts indicate inner and outer circles. According to Eq. (134), we get
$\mathcal{A}^{ii}_{\Gamma}+\mathcal{A}^{oo}_{\Gamma}=\frac{8}{3}\,\frac{R_{1}^{3}+R_{2}^{3}}{|\Gamma|^{2}}\,$, (147)
with $|\Gamma|=\pi(R_{2}^{2}-R_{1}^{2})$. We now compute the remaining contributions.
For the outer circle, we employ the parameterization $\bm{n}_{\bm{x}_{1}}=(\cos t_{1},\sin t_{1})^{\dagger}$, $\bm{x}_{1}(t_{1})=R_{2}\bm{n}_{\bm{x}_{1}}$, $dl_{\bm{x}_{1}}=R_{2}dt_{1}$, and similar for $\bm{x}_{2}$. For the inner circle, we have $\bm{n}_{\bm{x}_{1}}=-(\cos t_{1},\sin t_{1})^{\dagger}$, $\bm{x}_{1}(t_{1})=-R_{1}\bm{n}_{\bm{x}_{1}}$, $dl_{\bm{x}_{1}}=R_{1}dt_{1}$, and similar for $\bm{x}_{2}$. We get then
$\mathcal{A}^{io}_{\Gamma}$ $=\frac{R_{1}R_{2}\sqrt{R_{1}^{2}+R_{2}^{2}}}{3|\Gamma|^{2}}\int\limits_{0}^{2\pi}dt\biggl(\cos(t)\sqrt{1-\gamma\cos(t)}$
$+\frac{\cos(t)-\tfrac{\gamma}{2}(1+\cos^{2}(t))}{\sqrt{1-\gamma\cos(t)}}\biggr),$
where $\gamma=2R_{1}R_{2}/(R_{1}^{2}+R_{2}^{2})<1$. The last integral can be expressed in terms of the complete elliptic integrals as
$\mathcal{A}^{io}_{\Gamma}$ $=\frac{4R_{1}R_{2}\sqrt{R_{1}^{2}+R_{2}^{2}}}{3|\Gamma|^{2}\gamma\sqrt{1+\gamma}}\biggl((1-\gamma^{2})K(\sqrt{2\gamma/(1+\gamma)})$
$-(1+\gamma)E(\sqrt{2\gamma/(1+\gamma)})\biggr).$ (148)
By symmetry, one also gets $\mathcal{A}^{oi}_{\Gamma}=\mathcal{A}^{io}_{\Gamma}$ that completes the computation of $\mathcal{A}_{\Gamma}$.
For a thin annulus, one can set $R_{1}=R$ and $R_{2}=R+a$. Using the asymptotic behavior of the complete elliptic integrals $K(z)$ and $E(z)$ as $z\to 1$, we get after simplifications the small-$a$ behavior of $\mathcal{A}_{\Gamma}$:
$\mathcal{A}_{\Gamma}\approx\frac{\ln(8R/a)+3/2}{2\pi^{2}R}\,.$ (149)
Comparing this expression with Eq. (141) for a thin rectangle, we conjecture the following leading-order dependence for any thin patch, which can be seen as a “thickening” of a planar curve $C$ by “width” $2a$ (i.e., $\Gamma=\{\bm{x}\in\partial{\mathbb{R}}^{3}_{+}~:~|\bm{x}-C|<a\}$):
$\mathcal{A}_{\Gamma}\simeq\frac{2\ln(1/a)}{\pi|\partial\Gamma|}+O(1)\quad(a\to 0),$ (150)
where $|\partial\Gamma|$ is the perimeter of the patch. Combining this relation with the bounds (86) and assuming that $F_{0}$ is close to $1$, we get another conjecture that
$\frac{1}{a\mu_{0}}\simeq\frac{2}{\pi}\ln(1/a)+O(1)\,$, (151)
independently of the actual shape and length of the thin patch, which affect the constant (subleading) term.